### v2

# Randomly Sampled Language Reasoning Problems Reveal Limits of LLMs

Abstract
Can LLMs pick up language structure from examples? Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures and vocabularies. However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with these seen languages. In contrast, in this work we attempt to measure models’ language understanding capacity while circumventing the risk of dataset recall. We parameterize large families of language tasks recognized by deterministic finite automata (DFAs), and can thus sample novel language reasoning problems to fairly evaulate LLMs regardless of training data. We find that, even in the strikingly simple setting of 3-state DFAs, LLMs underperform unparameterized $n$-Gram models on both language recognition and synthesis tasks. These results suggest that LLMs struggle to match the ability of basic language models in recognizing and reasoning over languages that are sufficiently distinct from the ones they see at training time, underscoring the distinction between learning individual languages and possessing a general theory of language.
Machine Learning, ICML
1 Introduction
Figure 1: We use (a) randomly generated deterministic finite automata (DFAs) to sample previously unseen grammars to test models with. The DFA shown here, modeling the sum modulo 3 operation (with abc representing 0, 1, and 2 respectively), can be used to accept or reject strings from a 3-character alphabet. Accepted strings belong to the grammar, and rejected strings do not. We evaluate models on their ability to (b) act as a transducer, recognizing strings that belong to the DFA-defined grammar, and (c) generate new strings following the grammar.
Contemporary LLMs have proven themselves to be highly sophisticated natural language completion models that demonstrate many properties of reasoning engines. This has prompted questions surrounding the true intelligence of these models, with some arguing that they possess inherent language learning capabilities (Millière, 2024). In this paper, we explore the question of whether LLMs have the reasoning capacity to understand the structure of a new language. Specifically, we are interested in problems where a model is given a small set of examples from a language and either generates a new sample or determines whether a new sample is from the language or not.
Some work suggests that LLMs broadly understand language structure because they are able to produce syntactically correct samples from languages they have not been trained on (Athiwaratkun et al., 2022), although LLMs’ performance on low-resource languages tends to be lower than their ability on higher resource languages (Bogin et al., 2023). On the other hand, some critics of LLMs argue that LLMs cannot possess an understanding of language structure as they have learned from data rather than possessing a priori universal grammar (Chomsky et al., 2023). We do not view learning from data as a fundamental limitation, but we are concerned with the possibility that a language model might only be able to understand linguistic structures similar to those it has seen in training data. To distinguish between these possibilities, we wish to evaluate LLMs on a benchmark that considers wholly novel languages, eliminating the possibility of dataset leakage.
Ideally, models would be tested on a set of language reasoning problems disjoint from data seen during training and validation. However, as training datasets for LLMs are generally closed and incredibly vast, human-generated problems in natural language are likely to at least partially overlap in syntax or concept with content LLMs have already seen, making probing a model for its ability to reason about the structure of natural languages nearly impossible. Additionally, the problem of determining whether two tasks are semantically identical is itself a nontrivial one. Therefore, ensuring that even an entirely novel invented problem is not a variation on a theme is intractable.
To circumvent this problem, we propose the following general approach: first define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level. Then, sample random languages from this space. By sampling randomly from this space, we can guarantee no bias towards canonical languages that might share structure with ones in the training dataset. In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level, but this technique can be generalized to produce benchmarks of any difficulty level.
Experimental results using this approach suggest that contemporary LLMs possess less sophisticated language pattern recognition abilities than expected; underperforming basic, parameter-free n-gram language models on even the simplest languages. These results, combined with LLMs’ impressive results on a variety of specific tasks, suggest that LLMs function as ensemble models over tasks they have seen in their dataset, but do not possess the ability to generalize to entirely novel language reasoning tasks.
In summary, we make the following contributions:
  1. 1.
    We introduce a benchmark for LLM language reasoning evaluation, disjoint from natural language web data.
  2. 2.
    We evaluate a suite of popular LLMs on instances of this benchmark and demonstrate that LLMs underperform compared to simple language model baselines.
  3. 3.
    We analyze the differences in behavior between these models, illustrating the influence of RLHF and chain-of-thought prompting on language reasoning capacity.
2 Related Work
2.1 Reasoning with LLMs
Reasoning is one of many “emergent abilities” (Wei et al., 2022a) possibly possessed by LLMs (Huang & Chang, 2022), although the nonlinear dependence of such emergent abilities on model size is disputed (Schaeffer et al., 2024). The chain-of-thought prompting technique (Wei et al., 2022b) has inspired a number of approaches to encourage latent reasoning ability of models (Yao et al., 2023; Besta et al., 2024; Kojima et al., 2022), including neuro-symbolic methods (Hua & Zhang, 2022; Weir et al., 2023, 2024). Building on this, other work considers how to optimize exemplars used for in-context learning (Dong et al., 2022) and chain-of-thought prompting, known as “rationale refinement” (Liu et al., 2021; Fu et al., 2022). Problem-decomposition is also shown to be effective (Zhou et al., 2022; Khot et al., 2022).
2.2 LLM reasoning evaluation
LLM reasoning abilities are often tested on natural language benchmarks and commonly seen problems like arithmetic (Cobbe et al., 2021; Amini et al., 2019; Hendrycks et al., 2021), commonsense reasoning (Bhargava & Ng, 2022), and other, sometimes generative, tasks (Lake & Baroni, 2018; Pasupat & Liang, 2015; Lin et al., 2019) and task collections (Srivastava et al., 2022). LLMs have been shown to lack sufficient reasoning capability across a range of tasks including multi-step planning and complex inference (Valmeekam et al., 2022). Fan et al. (2023) introduce an LLM reasoning benchmark on algorithmic problems through NP-hard complexity, and Hazra et al. (2024) show that LLMs struggle to complete simple 3SAT problems. Patel et al. (2021) demonstrate that much of LLM mathematical reasoning can be explained by shallow heuristics, and Razeghi et al. (2022) similarly find that term frequency in training data impacts models’ in-context learning ability. In comparison to these, we attempt to explore the distinction described by Patel et al. (2021), but push both language simplicity and language unfamiliarity to their limits, by exploring simple languages recognized by randomly sampled DFAs. This enables us to evaluate the ability of LLMs to reason about language.
2.3 Language Understanding and LLMs
LLMs can be quite adept at generating programs in general-purpose programming languages (Xu et al., 2022a). In contrast, adapting models to understand domain-specific languages (Mernik et al., 2005) introduces unique problems such as navigating idiosyncratic syntax and semantics, and leveraging sparse collections of sample language data. To address these challenges, researchers have considered how well general-purpose LLMs can use language reasoning skills to quickly understand rare or unseen DSLs with only a small set of exemplars (Joel et al., 2024). While most work in this vein focuses on semantic parsing for downstream applications (Lin et al., 2023), selecting exemplars (Zhao et al., 2021), and improving DSL recognition by leveraging more common languages (Bogin et al., 2023), experiments show strong baseline performance for LLM DSL recognition and parsing out-of-the-box (Wang et al., 2024), indicating that LLMs may possess emergent language reasoning abilities.
Related lines of work are compositional generalization (Xu et al., 2022b), which assesses models’ ability to organize known units into novel structures, and structural generalization (Yao & Koller, 2022), which assesses models’ ability to recognize new structures. Yao & Koller (2022) show that smaller language models like BART and T5 can struggle on these tasks, but to our knowledge there are not comprehensive experiments extending this line of work to LLMs.
3 DFA Reasoning Tasks
3.1 DFAs and Regular Languages
The original Chomsky Hierarchy (Chomsky, 1959) separates language into four types (Figure 2). We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automaton (DFA). Examples of languages recognized by DFAs include simple examples like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7. Finite automaton languages are equivalent to regular languages, which are recognized by regular expressions.
One simple metric of the difficulty of a regular language is the number of states in the corresponding DFA, which represents the amount of memory the automaton has at any point while processing a given sequence. [^1] There are other metrics of difficulty, but we choose number of states as it is highly parsimonious. DFAs with 2 states have the property that their set of states is isomorphic to the output set $\{0,1\}$, and, therefore, do not have any hidden state. We thus explore 3-state DFAs, as this is the simplest nontrivial case.
Figure 2: An illustration of Chomsky’s hierarchy of languages, ranging from Type 0 to Type 3, which are defined by what formal models can recognize their grammars. In this work, we focus on the simplest language type in the hierarchy, regular grammars, which are recognized by deterministic finite automata (DFAs).
3.2 Sequence Completion Task
We also pose a sequence completion task, in which models must complete a sequence in a given DFA’s language. We study this task primarily because, in practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
To generate test cases for this task, we (1) sample a DFA, (2) sample 30 example sequences of length 10 that this DFA accepts, and (3) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept. The task is to find a completion [^2] Not necessarily the same completion found in sampling. of this prefix of between 1 and 5 characters such that the DFA accepts the full sequence. For details on sampling, see Appendix A.2.
We evaluate models by (1) sampling a DFA, (2) sampling 30 problem instances at random, and then (3) computing a binary prediction score (whether or not the predicted completion creates a valid string in the language) for each instance separately. We then average this metric over several sampled DFAs to produce a single accuracy score.
3.3 Transducer Task
While the sequence completion task is the natural one that comes to mind as a basic language task, it has a difficulty-gap problem. Specifically, the issue is that many DFAs, including the one shown in Figure 1, recognize languages that are particularly difficult to identify based on a set of examples, unless you build some kind of world model. Other DFAs end up being trivial to generate a completion for by analyzing common suffixes.333 The difficulty gap exists because a set of recognized sequences of length 10 gives no direct insight into intermediate states between the first and tenth token. As such, to be able to utilize this information for languages like the one in Figure 1 where there are no “resets” (sequences of symbols that necessarily lead to a particular state), a model must be capable of hollistically evaluating the entire sequence, probably requiring a world model. Many other DFAs contain these resets, but do so in such a way that makes it possible to e.g., recognize that all sequences in this language end in a, making the problem trivial. To provide a more direct evaluation of non-world-modeling-based pattern recognition, we explore the Transducer task.
In this task, an input sequence is annotated with an output at each token, the final output is masked, and the masked output is predicted by a language model. We call this a transducer task, as the DFA is used as a machine that converts a sequence of inputs into a sequence of outputs. E.g., given the language does the string have an even number of ’a’ tokens and the input abcabcaabbccaa, the annotated string (all that is provided to the model) is a0b0c0a1b1c1a0a1b1b1c1c1a0a and the output to predict is 1. For each problem instance, we provide 30 symbols, and for the first 29, the corresponding transducer output.
This task is significantly more transparent than the sequence completion task as the model has access to intermediate transducer outputs, an (imperfect) proxy for intermediate state. n
3.4 Baselines
l baseline models with varying degrees of sophistication.
Sequence Completion Task
For the Sequence Completion task, we have three kinds of baseline.
  • RandomS baseline: produce a random string of length 5 characters. While this might seem redundant as it should have a success rate of 50%, in practice our rejection sampling approach (see Appendix A.2) leads to a slight bias towards DFAs with more accept states. This baseline measures that bias.
  • Common-SuffixS baseline: find the completion $s$ of length between 1 and 5 that maximizes ((((# of occurrences as a ${}\times|s|)$. This baseline does not take the distinct prefix into account, and instead tries to find a universal completion that will always end in an accept state for this language.
  • $n$-GramS baseline: we take the last $n-1$ characters of the distinct prefix and search to see if they appear in any of the example sequences at a position where the sequence following is an appropriate length to be a completion (at least 1 but at most 5). We then take a plurality vote among the completions and return this, breaking ties arbitrarily. If there are no matches, we return the result of $(n-1)$-GramS. Technically these cover more than $n$ characters, since the completion is often $>1$ character long; for simplicity, however, we keep the naming consistent with the Transducer baselines.
  • Brute-ForceS: take all possible DFAs with 3 states and 3 symbols. Filter for ones that accept all the example sequences. Then try all remaining DFAs on all $3^{5}$ possible 5-length completions and return the completion that the maximal number of DFAs accept.
Note that these baselines are entirely unparameterized and operate identically regardless of the underlying DFA. This makes them direct comparisons to using LLMs in in-context-learning. We do not consider BruteForceS to be a reasonable comparison due to its computational complexity, and instead consider it an upper bound on performance on this particular task.
Transducer Task
We have similar baselines for the Transducer task.
  • NullT baseline: for a given DFA, whichever of the following strategies produces a better result: always predict 0 or always predict 1.
  • $n$-GramT baseline: take the $n-1$ symbols ending at the end of the concatenated transducer sequence (e.g., for $n=5$ and the above example, this would be 1a0a). If that sequence does not appear elsewhere in the sequence, return the result of the $(n-1)$-GramT baseline. Otherwise, take the token that appears immediately after each occurrence. If there is a majority, return that, otherwise return the last example.
  • BruteForceT: take all possible DFAs with 3 states and 3 symbols. Filter them for ones that match the given transducer sequence. Take this set and predict the next token. Take a majority vote among these, returning 1 by default if there is no majority.
4 Experiments
We evaluated the open-source models Llama 3-8B, Llama 3-70B (AI@Meta, 2023), and Llama 3-8B-Instruct (AI@Meta, 2024), Mistral Nemo Minitron 8B (NVIDIA, 2024), Mistral Nemo Base 2407 (Mistral AI, 2024b) and Mistral Nemo Instruct 2407 (Mistral AI, 2024c), Gemma 7B (Google, 2024), and Falcon 7B (Almazrouei et al., 2023).
We also evaluated the open-source code models StarCoder2-15B (Lozhkov et al., 2024), Codestral-22B-v0.1 (Mistral AI, 2024a), Deepseek Coder 33B Instruct (Deepseek, 2024), Qwen2.5-Coder-7B, Qwen2.5-Coder-7B-Instruct, and Qwen2.5-Coder-32B-Instruct (Hui et al., 2024).
Finally, we evaluated the proprietary models GPT-3.5-turbo-instruct, GPT-3.5 Chat (turbo-0125) (OpenAI, 2024a), GPT-4o-mini (2024-07-18), GPT 4o (2024-05-13) (OpenAI, 2024b), and Claude 3.5 Sonnet (Anthropic, 2024).
For each open source model, we used a local VLLM (Kwon et al., 2023) server for evaluation and always evaluated on 1000 distinct DFAs. For GPT-4o and Claude, we evaluated on 30 DFAs due to computation costs. For all other proprietary models, we evaluated on 100 DFAs. All results are reported with 95% bootstrap confidence intervals.
For both tasks, we consider four prompting formats:
  • Basic provides no context, presenting the problem as a generic next-token prediction or sequence generation task, where output is provided immediately following the input, with no space to think.
  • More-Expl provides some general information about the grammar that the input strings are generated from. This remains a next token prediction/sequence generation task.
  • COT provides general information about the grammar, and additionally invokes chain-of-thought reasoning to help the model reason over the task. Here, the model is given space to reason before providing a tagged answer.
  • Red-Green casts the tasks as independent word problems that describe the underlying grammar structure without relying on world knowledge about DFAs and regular languages. It describes an N-state DFA as a house with N rooms, each of which has 3 portals that deterministically go to other rooms (or back to the same room), where the walls of each room are red or green (mirroring transducer output symbols 0 and 1). Similarly to COT, the model is given space to show work before providing a tagged answer.
We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts. Full listings of these prompts can be found in Appendix D, with Table 4 containing a summary of each prompt. While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe our set of prompts captures common prompting strategies.
Model Size IT? Code? Sequence Completion SR Transducer TR
Baselines
BruteForce 100.0 (99.9–100.0) 1 96.4 (96.2–96.7) 1
6-Gram 91.7 (91.0–92.4) 2 93.5 (93.1–93.9) 2
5-Gram 91.2 (90.4–91.9) 3 93.4 (93.0–93.7) 3
4-Gram 89.6 (88.7–90.4) 4 91.1 (90.6–91.6) 4
3-Gram 87.0 (86.1–87.8) 5 87.0 (86.4–87.6) 16
2-Gram 83.3 (82.2–84.2) 8 74.5 (73.6–75.3) 24
Common-Suffix 84.7 (83.6–85.6) 6
Null 68.9 (68.2–69.6) 25
Random 53.3 (51.7–54.7) 26
Open Source Completion
llama3-8B 8.0B 73.8 (72.4–75.1) 18 87.5 (86.9–88.0) 14
llama3-70B 70.6B 71.4 (70.0–72.7) 23 87.7 (87.2–88.3) 12
llama3.1-8B-Instruct 8.0B 75.3 (74.0–76.6) 16 85.9 (85.3–86.5) 18
mistral-nemo-minitron-8B 8.4B 78.7 (77.5–79.8) 12 88.6 (88.0–89.1) 5
mistral-nemo-base-12B 12.2B 75.5 (74.3–76.6) 15 87.9 (87.4–88.4) 10
mistral-nemo-instruct-12B 12.2B 72.2 (70.9–73.4) 22 88.0 (87.5–88.5) 8
gemma-7b 8.5B 72.6 (71.3–73.7) 20 82.1 (81.4–82.7) 22
falcon-7b 7.2B 69.0 (67.6–70.2) 24 84.9 (84.3–85.5) 20
Open Source Code
starcoder2-15b 16.0B 73.5 (72.0–74.7) 19 87.7 (85.8–89.5) 13
codestral-22B 22.2B 78.0 (76.8–79.1) 13 86.6 (86.0–87.1) 17
deepseek-coder-33b-instruct 33.3B 76.7 (75.3–77.8) 14 85.6 (85.0–86.2) 19
qwen-2.5-coder-7B 7.6B 79.5 (78.4–80.5) 9 88.2 (87.6–88.7) 7
qwen-2.5-coder-instruct-7B 7.6B 79.5 (78.3–80.5) 10 88.3 (87.8–88.8) 6
qwen-2.5-coder-instruct-32B 32.8B 79.2 (78.0–80.3) 11 87.9 (87.4–88.4) 9
Proprietary
gpt-3.5-instruct ? 67.3 (63.1–71.5) 25 87.8 (85.9–89.6) 11
gpt-3.5-chat ? N/A 66.8 (63.4–69.8) 26
gpt-4o-mini ? 72.4 (68.1–76.3) 21 79.8 (77.3–82.2) 23
gpt-4o ? 74.4 (69.9–78.6) 17 83.7 (80.1–86.9) 21
claude-3.5 ? 84.0 (79.3–88.4) 7 87.1 (83.9–90.2) 15
Table 1: Results for our experiments. We present model metadata alongside model results on both the Transducer and Sequence completion tasks. Each cell contains the mean performance across DFAs for the best-performing prompt (see Table 2 for details), with 95% confidence intervals of the mean in parentheses. “N/A” is used whenever the model returned an invalid result at least 25% of the time. (IT = Instruction-Tuned, TR/SR = Transducer/Sequence Completion rank, the ordinal rank of the given model on the given task.)
5 Results
Main results for all tasks are presented in Table 1. For all LLMs, we ignore non-answers, i.e., if for a given DFA a model gets 25 correct answers, 1 incorrect answer, and responds with an unparseable result on 4, this counts as a 25/26, not a 25/29.
5.1 Sequence Completion
As seen in Table 1, this task is nearly always fully determined, that is, it can be solved with $\sim$100accuracy in theory, as demonstrated by BruteForceS results. Of course, BruteForceS is extremely computationally expensive, and, as such, we primarily focus on the $n$-GramS heuristics as our baselines. Still, we find, similarly to the transducer task, that $n$-GramS heuristics tend to outperform LLMs.
As seen in Table 2, we find that giving the model the opportunity to logically reason about the prompt via chain-of-thought and present a conclusion has inconsistent results. Specifically, we find that BasicS is the best prompt for gpt-4o-mini, but not gpt-4o, where the best performing prompt is Red-GreenS. We find that claude-3.5 is entirely unable to follow the sequence completion prompts BasicS and More-ExplS, and performs best at the COTS prompt.
Additionally, we find that in this task, code-specific open-source models tend to perform better than sequence completion models, suggesting some generalized ability to produce strings from novel languages demonstrated by example. Overall, the relative performances of LLMs and prompts comport somewhat well to heuristics on which models and prompting strategies should work best. Nonetheless, LLMs underperform simple $n$-Gram heuristics.
One potential problem with using this task for cross-model comparisons is the relevance of tokenization. Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix B for details.
Model Basic More-Expl COT Red-Green
Sequence Completion
gpt-4o-mini 72.4 (68.1–76.3) 70.5 (66.4–74.6) 58.0 (53.4–62.4) 59.1 (54.9–63.2)
gpt-4o 72.1 (65.9–78.2) N/A 67.4 (60.8–73.8) 74.4 (69.9–78.6)
claude-3.5 N/A N/A 84.0 (79.3–88.4) 80.0 (74.9–85.2)
Transducer
gpt-4o-mini 79.8 (77.3–82.2) 76.7 (74.2–79.3) 65.2 (63.1–67.4) 74.5 (72.0–77.0)
gpt-4o 83.7 (80.1–86.9) 82.6 (79.1–85.9) 67.8 (63.1–72.3) 82.6 (78.8–86.3)
claude-3.5 86.9 (83.3–90.0) 87.1 (83.9–90.2) 76.4 (72.9–79.9) 82.9 (78.9–86.9)
Table 2: Results for chat models (where we investigated multiple prompts). For all other models, we only used the Basic prompt. We find that non-COT prompts consistently work better for the Transducer task, with more mixed results on sequence completion.
5.2 Transducer
Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ model only achieving 96.4% accuracy. However, relative results should still be valid as the impossible problems are present with equal probability for all models.
We find that in general all LLMs underperform a 4-GramT model, demonstrating that they are unable to adequately solve this task. The relative performance of the models also does not generally correspond to their overall scale, with open source LLama-3 and Mistral Nemo 8B parameter models outperforming Claude and GPT-4o. Even within a model class we find no clear pattern: GPT-4o is outperformed by GPT 3.5, Llama 3-70B has similar performance to Llama 3-8B, and the Mistral Nemo 12B models perform similarly to Nemo Minitron 8B. Coding models also demonstrate no particular advantage on this task.
The generally lower performance of chat-oriented models suggests this task is better suited to non-chat models. To investigate that this is not specific to the Basic prompt, we investigate several chat models on the other prompts. As seen in Table 2, our chain-of-though and word problem prompts, which attempt to leverage the full reasoning capabilities of chat models, also fare poorly, performing similarly or worse to the Basic prompt on the Transducer task.
Overall, we conclude that LLMs are unable to perform the DFA transducer inference task to a reasonable degree. This failure cannot be attributed to a lack of world modeling ability, as $n$-GramT models do not construct world models. Instead, it seems the LLMs are unable to detect patterns when those patterns are drawn from an unfamiliar source, even a relatively simple one.
5.3 Case Study: Sum Modulo 3 DFA
Correct Incorrect
Total 30 30
a is no-op 21 22
1b and 1c lead to 0 14 17
2 red rooms 2 3
2-periodic 9 14
3-periodic 4 4
$>3$-Gram 4 5
Sum Modulo 3 0 0
Figure 3: Results on Sum Modulo 3 DFA. (a) Venn diagram results for CR (claude-3.5 using Red-GreenT), QB (qwen-2.5-coder-instruct-7B using BasicT), and 6-GramT. Percentages are accuracy numbers, and venn diagram is error counts, out of 1000. (b) Results of qualitative analysis. 30 CR reasoning traces were sampled in each condition, with Incorrect traces guaranteed to correspond to problems 6-GramT got correct.
We focus on the sum modulo 3 DFA in this section. We depict this DFA in Figure 1. This DFA can be interpreted as an arithmetic check, where a represents 0, b represents 1, and c represents 2, and the DFA accepts strings whose sum is equal to 0 modulo 3. We focus on the Transducer setting as sum modulo 3 is nearly impossible (for LLMs and $n$-Gram models alike) in the sequence completion setting due to the issues raised in Section 3.3. For this case study, we focus on the model/prompt combinations QB and CR, defined as
  • QB: qwen-2.5-coder-instruct-7B with BasicT. Selected as it is the best performing combination overall.
  • CR: claude-3.5 with Red-GreenT. Selected as it is the best performing combination that provides an explanation (needed later for our qualitative analysis)
Figure 3a depicts the number of errors each model receives on 1000 instances of the transducer task for this DFA. In general, nearly all errors made by the 6-GramT model were also made by at least one LLM model, while the two LLM models often made unique errors. Interestingly, while this task is better-known than most DFAs, we find that all 3 models perform worse on this DFA than their average across random DFAs.
We also performed a qualitative analysis, investigating CR’s outputs on the Red-GreenT prompt to see what kind of reasoning it is using; specifically we sampled 30 examples where it had the correct answer, and 30 examples where it had the incorrect answer but the 6-GramT model had the correct answer. Results of this analysis can be found in Figure 3b. We find that, in general, CR is following a 3-Gram approach, learning rules relating to the conditions under which the previous output and symbol can be used to predict the next output. Specifically, it is able to learn that a does not change the output, and that b and c will lead a 1 state to a 0 state. These results comport with the overall finding of Table 1, where we found that 3-GramT was the largest $n$-GramT that any LLM outperformed.
[…]
Despite transformers’ theoretically high computational capacity, without the ability to pattern match to existing problems, Claude uses an unsophisticated and ineffectual approach.
6 Conclusion
Our findings highlight significant weaknesses in large language models’ ability to generalize to entirely novel language reasoning problems, even simple ones solely involving next-token prediction on basic languages recognized by 3-state DFAs. These results, combined with that of previous work demonstrating that large language models can quite accurately perform a variety of language tasks, suggests that LLMs solve language problems via a mechanism distinct from general language reasoning ability. Our use of n-gram baselines and next-token prediction tasks allows us to exclude the possibility that the issue is primarily related to LLMs’ lack of world modeling or any inherent limitations of next-token prediction models. We believe our results suggest that LLMs have learned individual models of particular languages, but not a general theory of language.
Interestingly, in our transducer experiments, LLMs consistently perform better by directly predicting the next token than by explicitly reasoning through the problem. While our conclusions are limited by the finite nature of our prompt set, this suggests that they do, in fact, possess some latent understanding of language, but this understanding is inferior to basic n-gram models for $n>3$.
These results expose a particular weakness in LLMs, as many potential foundation model applications involve tasks that are not expressed in familiar human languages or pre-existing programming languages. More specifically, in tasks where there is a need to produce an output in a precise, atypical, format, we should be skeptical of the ability of LLMs to in-context-learn this format. For these tasks, it may be prudent to seek a new approach.
Acknowledgments
We thank Marc Marone and Theo Olausson for their advice on experiment design.
Appendix A Details on Sampling
A.1 Sampling of DFAs
We use rejection sampling to sample DFAs. Specifically, we uniformly sample a start state, then for each (source state, symbol) pair, we sample a post-transition state. We also randomly assign each state to be accept or reject with probability 50%. We then reject any DFA that has all accept or all reject states, or for which certain states are unreachable from the start state.
A.2 Sampling of Sequence Completion Tasks
To sample a sequence completion task, we first sample a DFA as described in Appendix A.1.
To sample a task instance, we sample example sequences and distinct prefix. Each example sequence is sampled uniformly from the space of $\{a,b,c\}^{10}$ and then rejected if the DFA does not accept the sequence. Our distinct prefix and completion are sampled uniformly from $\{a,b,c\}^{5}\times\{a,b,c\}^{5}$, and is rejected if the DFA does not accept the concatenation of the two, or if the prefix is the prefix of any of the previous sequences. We then discard the completion. If we, at any point, reject 50 sequences when attempting to sample a sequence or prefix, we return an error.
We run a “pilot” sampling for a DFA to ensure that it is valid, in which we sample an instance as described abov. If there is an error in sampling this pilot instance, we reject the DFA. Otherwise, we proceed to sample our task instances. At this stage, if there is an error in sampling, we reject the instance rather than the DFA.
A.3 Sampling of Transducer Tasks
We sample a DFA as described in Appendix A.1, and then sample random sequences (30 in our experiments) and generate transducer traces. If every transducer trace ends with a 0 or every trace ends with a 1, we reject the DFA and resample.
Model BasicS Basic-CommasS
qwen-2.5-coder-7B 79.5 (78.4–80.5) 60.7 (59.3–62.1)
qwen-2.5-coder-instruct-7B 79.5 (78.3–80.5) 55.5 (54.0–56.9)
qwen-2.5-coder-instruct-32B 79.2 (78.0–80.3) 55.2 (53.7–56.7)
mistral-nemo-minitron-8B 78.7 (77.5–79.8) 59.3 (57.9–60.8)
codestral-22B 78.0 (76.8–79.1) 59.0 (57.5–60.3)
deepseek-coder-33b-instruct 76.7 (75.3–77.8) 54.9 (53.0–56.8)
mistral-nemo-base-12B 75.5 (74.3–76.6) 60.6 (59.1–62.2)
llama3.1-8B-Instruct 75.3 (74.0–76.6) 56.3 (54.4–58.1)
llama3-8B 73.8 (72.4–75.1) 61.5 (60.2–62.9)
starcoder2-15b 73.5 (72.0–74.7) 58.2 (56.7–59.8)
gemma-7b 72.6 (71.3–73.7) 54.0 (51.9–56.0)
gpt-4o-mini 72.4 (68.1–76.3) 64.1 (59.5–68.3)
mistral-nemo-instruct-12B 72.2 (70.9–73.4) 58.2 (56.4–59.8)
gpt-4o 72.1 (65.9–78.2) 66.8 (58.5–74.8)
llama3-70B 71.4 (70.0–72.7) 56.4 (54.7–58.0)
falcon-7b 69.0 (67.6–70.2) 56.1 (54.5–57.6)
gpt-3.5-instruct 67.3 (63.1–71.5) 52.3 (46.5–57.9)
claude-3.5 N/A N/A
gpt-3.5-chat N/A N/A
Table 3: Results on Sequence Completion Task. We compare BasicS to the comma-variant Basic-CommasS.
Appendix B Sequence Completion task prompt with Commas
To avoid tokenization differences with models, we also investigate a version of our Sequence Completion prompt that uses spaces and commas between the elements of the sequence. Unfortunately, results using this prompt were uniformly worse than results on the prompt without spaces and commas. Table 3 shows the results on a variety of models. All are worse with commas than without.
Appendix C More details on Sum Modulo 3 DFA case study
Figure 4: Results on Sum Modulo 3 DFA under trivial / nontrivial conditions. Percentages are accuracy numbers, and venn diagram is error counts. (a) In this condition, CR and the 6-GramT both get very high accuracies, with nearly all 6-GramT also being CR errors. QB does poorly. (b) In this condition, models do significantly more poorly overall, with CR in particular performing worse than chance. Here, errors are more symmetric, with more 6-GramT errors that are not accounted for by either or both model, indicating that a larger fraction of both successes and failures in this condition are down to random chance.
Appendix D Prompt Listings
Prompt $T$ $S$
Basic You are a sequence completion model. Output the next element of the sequence, and nothing else. <TRANSDUCER PREFIX>, The following strings come from an alien language that follows a simple grammar. Infer the alien grammar using the example strings. Then, add a suffix to the final string using between 1 and 5 characters such that the full string follows the grammar. Output only the necessary suffix to complete the final string, and nothing else. <EXAMPLES> <PREFIX>
More-Expl You are a sequence completion model. The following sequence is generated from an unknown but consistent grammar. Identify the patterns within the sequence to determine its next element. Output the next element of the sequence, and nothing else. <TRANSDUCER PREFIX>, I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect a set of valid strings using this DFA, listed below. Infer the underlying DFA model using these strings and complete the final string, using up to n characters, such that it is also a valid string. Output only the necessary suffix to complete the final string, and nothing else. <EXAMPLES> <PREFIX>
COT A DFA is a finite-state machine that accepts or rejects a given string of symbols, by running through a n-state sequence uniquely determined by the string. I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect the inputs and outputs into an input sequence and an output sequence. Infer the underlying DFA model to predict the next integer in the output sequence. Reason step by step, and then output the next output integer using <answer> tags, like <answer>0</answer>. Input sequence: <TRANSDUCER PREFIX> Output sequence: I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect a set of valid strings using this DFA, listed below. Infer the underlying DFA model using these strings and complete the final string, using up to n characters, such that it is also a valid string. Reason step by step, and then output the next necessary suffix for this final string, <answer> tags, like <answer>ab</answer>. Given these valid strings: <EXAMPLES> Complete the following string: <PREFIX>
Red-Green ‘‘‘ You are in a house of rooms and portals. There are 3 rooms in the house, and each room has 3 unique portals labeled A, B, and C. Each portal teleports you to one room of the house (and sometimes the destination is the room the portal is in). Every portal in a given room always behaves the same way. In this house, each of the rooms look exactly the same, except some of the rooms have red walls and some have green walls. However, there are *three* rooms in total, so you cannot determine which room you are in by color alone, and two rooms of the same color may have portals that behave differently. As you move through the house, at each time step you write down what portal you take and the color of the room you arrive (or stay) in. Based on your notes, predict what color room you will end up in after the last step. Tag your final answer like <answer>color</answer>. You walk through a portal labeled "<TRANSDUCER PREFIX>" and end up in a red room. ‘‘‘ You are outside a house of rooms and portals. There are 3 rooms in the house, and each room has 3 unique portals labeled a, b, and c. Each portal teleports you to one room of the house (and sometimes the destination is the room the portal is in). Every portal in a given room always behaves the same way. In this house, each of the rooms look exactly the same, except some of the rooms have red walls and some have green walls. However, there are *3* rooms in total, so you cannot determine which room you are in by color alone, and two rooms of the same color may have portals that behave differently. Youve been into this house many times before. Each time, as you move through the house, you write down what series of portals you take and the color of the room you end up in. You have a collection of paths youve taken where youve ended up in a room with green walls, listed below. Given the final incomplete path at the bottom, write a series of up to 5 remaining steps that will cause you to end up in a room with green walls again. Tag your final answer like <answer>ab</answer>. Given these paths that end in a room with green walls: <EXAMPLES> Complete the following path: <PREFIX>
Table 4: Shortened summary of each prompt
D.1 Summaries
Table 4 contains summaries of each prompt.
D.2 Full example listings
D.2.1 BasicT
You are a sequence completion model. Output the next element of the sequence, and nothing else.
a, 1, b, 1, a, 1, b, 1, b, 1, c, 0, a, 1, c, 1, a, 1, a, 1, a, 1, c, 1, b, 1, c, 0, c, 1, a, 1, b, 1, b, 1, b, 1, b, 1, a, 1, b, 1, a, 1, a, 1, b, 1, c, 0, a, 1, c, 1, a, 1, b,
D.2.2 More-ExplT
You are a sequence completion model. The following sequence is generated from an unknown but consistent grammar. Identify the patterns within the sequence to determine its next element. Output the next element of the sequence, and nothing else.
a, 1, b, 1, a, 1, b, 1, b, 1, c, 0, a, 1, c, 1, a, 1, a, 1, a, 1, c, 1, b, 1, c, 0, c, 1, a, 1, b, 1, b, 1, b, 1, b, 1, a, 1, b, 1, a, 1, a, 1, b, 1, c, 0, a, 1, c, 1, a, 1, b,
D.2.3 COTT
A DFA is a finite-state machine that accepts or rejects a given string of symbols, by running through a n-state sequence uniquely determined by the string.
I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect the inputs and outputs into an input sequence and an output sequence. Infer the underlying DFA model to predict the next integer in the output sequence. Reason step by step, and then output the next output integer using <answer> tags, like <answer>0</answer>.
Input sequence: a, b, a, b, b, c, a, c, a, a, a, c, b, c, c, a, b, b, b, b, a, b, a, a, b, c, a, c, a, b
Output sequence: 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1,
D.2.4 Red-GreenT
You are in a house of rooms and portals. There are 3 rooms in the house, and each room has 3 unique portals labeled A, B, and C. Each portal teleports you to one room of the house (and sometimes the destination is the room the portal is in). Every portal in a given room always behaves the same way.
In this house, each of the rooms look exactly the same, except some of the rooms have red walls and some have green walls. However, there are *three* rooms in total, so you cannot determine which room you are in by color alone, and two rooms of the same color may have portals that behave differently. As you move through the house, at each time step you write down what portal you take and the color of the room you arrive (or stay) in. Based on your notes, predict what color room you will end up in after the last step.
Tag your final answer like <answer>color</answer>.
You walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a red room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a red room.
Then, you walk through a portal labeled "C" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a red room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "C" and end up in a green room.
Then, you walk through a portal labeled "A" and end up in a green room.
Then, you walk through a portal labeled "B" and end up in a ...
D.2.5 BasicS
The following strings come from an alien language that follows a simple grammar. Infer the alien grammar using the example strings. Then, add a suffix to the final string using between 1 and 5 characters such that the full string follows the grammar. Output only the necessary suffix to complete the final string, and nothing else.
cbcbabbcca
abcaaacbaa
aabccbabbb
bbbccbbbca
aababaccba
aaaacbacac
baacbccbaa
cbbaacabcc
baabaacaab
bbbbbcacab
acaabcbbba
acaacbccac
cacbabcbba
abcbcbcbcc
ccaccccaba
bcbcabbcca
baabacabca
caababacac
bacacaccaa
bcacbbbbca
bcbbbcaccc
ccabbcccbb
bccbcabbca
baacbabcbc
ccacabccab
caacbcaaab
cacbaaccac
aaccbcaabb
abacabcaab
bacbcbcaca
caacb
D.2.6 Basic-CommasS
The following strings come from an alien language that follows a simple grammar. Infer the alien grammar using the example strings. Then, add a suffix to the final string using between 1 and 5 characters such that the full string follows the grammar. Output only the necessary suffix to complete the final string, and nothing else.
c, b, c, b, a, b, b, c, c, a
a, b, c, a, a, a, c, b, a, a
a, a, b, c, c, b, a, b, b, b
b, b, b, c, c, b, b, b, c, a
a, a, b, a, b, a, c, c, b, a
a, a, a, a, c, b, a, c, a, c
b, a, a, c, b, c, c, b, a, a
c, b, b, a, a, c, a, b, c, c
b, a, a, b, a, a, c, a, a, b
b, b, b, b, b, c, a, c, a, b
a, c, a, a, b, c, b, b, b, a
a, c, a, a, c, b, c, c, a, c
c, a, c, b, a, b, c, b, b, a
a, b, c, b, c, b, c, b, c, c
c, c, a, c, c, c, c, a, b, a
b, c, b, c, a, b, b, c, c, a
b, a, a, b, a, c, a, b, c, a
c, a, a, b, a, b, a, c, a, c
b, a, c, a, c, a, c, c, a, a
b, c, a, c, b, b, b, b, c, a
b, c, b, b, b, c, a, c, c, c
c, c, a, b, b, c, c, c, b, b
b, c, c, b, c, a, b, b, c, a
b, a, a, c, b, a, b, c, b, c
c, c, a, c, a, b, c, c, a, b
c, a, a, c, b, c, a, a, a, b
c, a, c, b, a, a, c, c, a, c
a, a, c, c, b, c, a, a, b, b
a, b, a, c, a, b, c, a, a, b
b, a, c, b, c, b, c, a, c, a
c, a, a, c, b,
D.2.7 More-ExplS
I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect a set of valid strings using this DFA, listed below. Infer the underlying DFA model using these strings and complete the final string, using up to n characters, such that it is also a valid string. Output only the necessary suffix to complete the final string, and nothing else.
cbcbabbcca
abcaaacbaa
aabccbabbb
bbbccbbbca
aababaccba
aaaacbacac
baacbccbaa
cbbaacabcc
baabaacaab
bbbbbcacab
acaabcbbba
acaacbccac
cacbabcbba
abcbcbcbcc
ccaccccaba
bcbcabbcca
baabacabca
caababacac
bacacaccaa
bcacbbbbca
bcbbbcaccc
ccabbcccbb
bccbcabbca
baacbabcbc
ccacabccab
caacbcaaab
cacbaaccac
aaccbcaabb
abacabcaab
bacbcbcaca
caacb
D.2.8 COTS
I have a 3-state DFA model that outputs either 0 or 1 after each element I input. 1 indicates that the input string thus far results in a "valid" state, and 0 indicates that it does not. I collect a set of valid strings using this DFA, listed below. Infer the underlying DFA model using these strings and complete the final string, using up to n characters, such that it is also a valid string. Reason step by step, and then output the next necessary suffix for this final string, <answer> tags, like <answer>ab</answer>.
Given these valid strings:
cbcbabbcca
abcaaacbaa
aabccbabbb
bbbccbbbca
aababaccba
aaaacbacac
baacbccbaa
cbbaacabcc
baabaacaab
bbbbbcacab
acaabcbbba
acaacbccac
cacbabcbba
abcbcbcbcc
ccaccccaba
bcbcabbcca
baabacabca
caababacac
bacacaccaa
bcacbbbbca
bcbbbcaccc
ccabbcccbb
bccbcabbca
baacbabcbc
ccacabccab
caacbcaaab
cacbaaccac
aaccbcaabb
abacabcaab
bacbcbcaca
Complete the following string:
caacb
D.2.9 Red-GreenS
You are outside a house of rooms and portals. There are 3 rooms in the house, and each room has 3 unique portals labeled a, b, and c. Each portal teleports you to one room of the house (and sometimes the destination is the room the portal is in). Every portal in a given room always behaves the same way.
In this house, each of the rooms look exactly the same, except some of the rooms have red walls and some have green walls. However, there are *3* rooms in total, so you cannot determine which room you are in by color alone, and two rooms of the same color may have portals that behave differently. Youve been into this house many times before. Each time, as you move through the house, you write down what series of portals you take and the color of the room you end up in. You have a collection of paths youve taken where youve ended up in a room with green walls, listed below. Given the final incomplete path at the bottom, write a series of up to 5 remaining steps that will cause you to end up in a room with green walls again.
Tag your final answer like <answer>ab</answer>.
Given these paths that end in a room with green walls:
cbcbabbcca
abcaaacbaa
aabccbabbb
bbbccbbbca
aababaccba
aaaacbacac
baacbccbaa
cbbaacabcc
baabaacaab
bbbbbcacab
acaabcbbba
acaacbccac
cacbabcbba
abcbcbcbcc
ccaccccaba
bcbcabbcca
baabacabca
caababacac
bacacaccaa
bcacbbbbca
bcbbbcaccc
ccabbcccbb
bccbcabbca
baacbabcbc
ccacabccab
caacbcaaab
cacbaaccac
aaccbcaabb
abacabcaab
bacbcbcaca
Complete the following path:
caacb

### v3

@@
 Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures and vocabularies.
-However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with these seen languages.
+However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with seen languages.
 In contrast, in this work we attempt to measure models’ language understanding capacity while circumventing the risk of dataset recall.
-We parameterize large families of language tasks recognized by deterministic finite automata (DFAs), and can thus sample novel language reasoning problems to fairly evaulate LLMs regardless of training data.
+We parameterize large families of language tasks recognized by deterministic finite automata (DFAs), and can thus sample novel language reasoning problems to fairly evaluate LLMs regardless of training data.
 We find that, even in the strikingly simple setting of 3-state DFAs, LLMs underperform unparameterized $n$-Gram models on both language recognition and synthesis tasks.
-These results suggest that LLMs struggle to match the ability of basic language models in recognizing and reasoning over languages that are sufficiently distinct from the ones they see at training time, underscoring the distinction between learning individual languages and possessing a general theory of language.
+These results suggest that LLMs struggle to match the ability of basic language models in recognizing and reasoning over languages that are sufficiently distinct from the ones seen at training time, underscoring the distinction between learning individual languages and possessing a general theory of language.
 Machine Learning, ICML
@@
 Then, sample random languages from this space.
-By sampling randomly from this space, we can guarantee no bias towards canonical languages that might share structure with ones in the training dataset.
+By sampling randomly, we can guarantee no bias towards canonical languages that might share structure with ones in the training dataset.
 In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level, but this technique can be generalized to produce benchmarks of any difficulty level.
 Experimental results using this approach suggest that contemporary LLMs possess less sophisticated language pattern recognition abilities than expected; underperforming basic, parameter-free n-gram language models on even the simplest languages.
-These results, combined with LLMs’ impressive results on a variety of specific tasks, suggest that LLMs function as ensemble models over tasks they have seen in their dataset, but do not possess the ability to generalize to entirely novel language reasoning tasks.
+These results, combined with LLMs’ impressive results on a variety of specific tasks, suggest that LLMs function as ensemble models over language tasks they have seen in their dataset, but do not possess the ability to generalize to entirely novel language reasoning tasks.
 In summary, we make the following contributions:
@@
 Reasoning is one of many “emergent abilities” (Wei et al., 2022a) possibly possessed by LLMs (Huang & Chang, 2022), although the nonlinear dependence of such emergent abilities on model size is disputed (Schaeffer et al., 2024).
-The chain-of-thought prompting technique (Wei et al., 2022b) has inspired a number of approaches to encourage latent reasoning ability of models (Yao et al., 2023; Besta et al., 2024; Kojima et al., 2022), including neuro-symbolic methods (Hua & Zhang, 2022; Weir et al., 2023, 2024).
+The chain-of-thought prompting technique (Wei et al., 2022b) has inspired a number of approaches to encourage the latent reasoning ability of models (Yao et al., 2023; Besta et al., 2024; Kojima et al., 2022), including neuro-symbolic methods (Hua & Zhang, 2022; Weir et al., 2023, 2024).
 Building on this, other work considers how to optimize exemplars used for in-context learning (Dong et al., 2022) and chain-of-thought prompting, known as “rationale refinement” (Liu et al., 2021; Fu et al., 2022).
@@
 (2022) similarly find that term frequency in training data impacts models’ in-context learning ability.
-In comparison to these, we attempt to explore the distinction described by Patel et al.
+In comparison to these, we explore the distinction described by Patel et al.
 (2021), but push both language simplicity and language unfamiliarity to their limits, by exploring simple languages recognized by randomly sampled DFAs.
@@
 We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automaton (DFA).
-Examples of languages recognized by DFAs include simple examples like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7.
+Examples of languages recognized by DFAs include simple ones like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7.
 Finite automaton languages are equivalent to regular languages, which are recognized by regular expressions.
@@
 [^1] There are other metrics of difficulty, but we choose number of states as it is highly parsimonious.
-DFAs with 2 states have the property that their set of states is isomorphic to the output set $\{0,1\}$, and, therefore, do not have any hidden state.
+DFAs with 2 states have the property that their set of states is no larger than the output set $\{0,1\}$, and, therefore, do not have any hidden state.
 We thus explore 3-state DFAs, as this is the simplest nontrivial case.
@@
 3.2 Sequence Completion Task
-We also pose a sequence completion task, in which models must complete a sequence in a given DFA’s language.
+We first pose a sequence completion task, in which models must complete a sequence in a given DFA’s language.
 We study this task primarily because, in practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
-To generate test cases for this task, we (1) sample a DFA, (2) sample 30 example sequences of length 10 that this DFA accepts, and (3) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept.
-The task is to find a completion [^2] Not necessarily the same completion found in sampling.
-of this prefix of between 1 and 5 characters such that the DFA accepts the full sequence.
+To generate test cases for this task given a DFA, we first (1) sample 30 example sequences of length 10 that this DFA accepts, and then (2) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept.
+The task is to find a completion (not necessarily the same completion found in sampling) of this prefix of between 1 and 5 characters such that the DFA accepts the full sequence.
 For details on sampling, see Appendix A.2.
-We evaluate models by (1) sampling a DFA, (2) sampling 30 problem instances at random, and then (3) computing a binary prediction score (whether or not the predicted completion creates a valid string in the language) for each instance separately.
-We then average this metric over several sampled DFAs to produce a single accuracy score.
+We evaluate models by (1) sampling a DFA, (2) sampling 30 problem instances at random (each of which contains 30 example sequences and a distinct prefix), and then (3) computing a binary prediction score (whether or not the predicted completion creates a valid string in the language) for each instance separately.
+We then average this metric over several sampled DFAs to produce our accuracy score.
 3.3 Transducer Task
@@
 Specifically, the issue is that many DFAs, including the one shown in Figure 1, recognize languages that are particularly difficult to identify based on a set of examples, unless you build some kind of world model.
-Other DFAs end up being trivial to generate a completion for by analyzing common suffixes.333 The difficulty gap exists because a set of recognized sequences of length 10 gives no direct insight into intermediate states between the first and tenth token.
+Other DFAs end up being trivial to generate a completion for by analyzing common suffixes.222 The difficulty gap exists because a set of recognized sequences of length 10 gives no direct insight into intermediate states between the first and tenth token.
 As such, to be able to utilize this information for languages like the one in Figure 1 where there are no “resets” (sequences of symbols that necessarily lead to a particular state), a model must be capable of hollistically evaluating the entire sequence, probably requiring a world model.
-Many other DFAs contain these resets, but do so in such a way that makes it possible to e.g., recognize that all sequences in this language end in a, making the problem trivial.
+Many other DFAs contain these resets, but do so in such a way that makes it possible to e.g., recognize that all sequences that end in a are in the language, making the problem trivial.
 To provide a more direct evaluation of non-world-modeling-based pattern recognition, we explore the Transducer task.
@@
 3.4 Baselines
-l baseline models with varying degrees of sophistication.
+l To contextualize LLM accuracies, waseline models with varying degrees of sophistication.
 Sequence Completion Task
@@
 Filter for ones that accept all the example sequences.
-Then try all remaining DFAs on all $3^{5}$ possible 5-length completions and return the completion that the maximal number of DFAs accept.
+Then try all remaining DFAs on all $3^{5}$ possible 5-length completions and return the completion that the maximal number of DFAs accept, breaking ties arbitrarily.
 Note that these baselines are entirely unparameterized and operate identically regardless of the underlying DFA.
@@
 We have similar baselines for the Transducer task.
-  • NullT baseline: for a given DFA, whichever of the following strategies produces a better result: always predict 0 or always predict 1.
+  • NullT baseline: for a given DFA, whichever of the following strategies produces a higher accuracy: always predict 0 or always predict 1.
   • $n$-GramT baseline: take the $n-1$ symbols ending at the end of the concatenated transducer sequence (e.g., for $n=5$ and the above example, this would be 1a0a).
@@
 We also evaluated the open-source code models StarCoder2-15B (Lozhkov et al., 2024), Codestral-22B-v0.1 (Mistral AI, 2024a), Deepseek Coder 33B Instruct (Deepseek, 2024), Qwen2.5-Coder-7B, Qwen2.5-Coder-7B-Instruct, and Qwen2.5-Coder-32B-Instruct (Hui et al., 2024).
-Finally, we evaluated the proprietary models GPT-3.5-turbo-instruct, GPT-3.5 Chat (turbo-0125) (OpenAI, 2024a), GPT-4o-mini (2024-07-18), GPT 4o (2024-05-13) (OpenAI, 2024b), and Claude 3.5 Sonnet (Anthropic, 2024).
+Finally, we evaluated the proprietary models GPT-3.5-turbo-instruct, GPT-3.5 Chat (turbo-0125) (OpenAI, 2024a), GPT-4o-mini (2024-07-18), GPT 4o (2024-05-13) (OpenAI, 2024b), o1-preview (2024-09-12) (OpenAI et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024).
 For each open source model, we used a local VLLM (Kwon et al., 2023) server for evaluation and always evaluated on 1000 distinct DFAs.
 For GPT-4o and Claude, we evaluated on 30 DFAs due to computation costs.
-For all other proprietary models, we evaluated on 100 DFAs.
-All results are reported with 95% bootstrap confidence intervals.
+For o1-preview we evaluated on only 10 DFAs, and only on the Transducer task (which we felt was a better fit for a reasoning model).
+For gpt-3.5 and gpt-4o-mini, we evaluated on 100 DFAs.
+All models were evaluated with temperature 0, except o1-preview333o1-preview does not allow setting a non-default temperature..
 For both tasks, we consider four prompting formats:
-  • Basic provides no context, presenting the problem as a generic next-token prediction or sequence generation task, where output is provided immediately following the input, with no space to think.
-  • More-Expl provides some general information about the grammar that the input strings are generated from. This remains a next token prediction/sequence generation task.
-  • COT provides general information about the grammar, and additionally invokes chain-of-thought reasoning to help the model reason over the task.
+  • Basic provides no context, presenting the problem as a generic sequence generation or next-token prediction task, where output is provided immediately following the input, with no space to think.
+  • More-Expl explains that the strings are generated from a DFA with 3 states, but is otherwise identical to Basic. This remains a sequence generation/next token prediction task.
+  • COT provides the same information as More-Expl and additionally invokes chain-of-thought reasoning to help the model reason over the task.
 Here, the model is given space to reason before providing a tagged answer.
@@
 We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts.
-Full listings of these prompts can be found in Appendix D, with Table 4 containing a summary of each prompt.
+Full listings of these prompts can be found in Appendix E, with Table 5 containing a summary of each prompt.
 While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe our set of prompts captures common prompting strategies.
@@
 3-Gram 87.0 (86.1–87.8) 5 87.0 (86.4–87.6) 16
-2-Gram 83.3 (82.2–84.2) 8 74.5 (73.6–75.3) 24
+2-Gram 83.3 (82.2–84.2) 8 74.5 (73.6–75.3) 25
 Common-Suffix 84.7 (83.6–85.6) 6
-Null 68.9 (68.2–69.6) 25
-Random 53.3 (51.7–54.7) 26
+RandomS/NullT 53.3 (51.7–54.7) 26 68.9 (68.2–69.6) 26
 Open Source Completion
@@
 gpt-3.5-instruct ? 67.3 (63.1–71.5) 25 87.8 (85.9–89.6) 11
-gpt-3.5-chat ? N/A 66.8 (63.4–69.8) 26
+gpt-3.5-chat ? N/A 66.8 (63.4–69.8) 27
 gpt-4o-mini ? 72.4 (68.1–76.3) 21 79.8 (77.3–82.2) 23
@@
 claude-3.5 ? 84.0 (79.3–88.4) 7 87.1 (83.9–90.2) 15
+o1-preview ? 76.5 (69.4–84.3) 24
 Table 1: Results for our experiments.
@@
 For all LLMs, we ignore non-answers, i.e., if for a given DFA a model gets 25 correct answers, 1 incorrect answer, and responds with an unparseable result on 4, this counts as a 25/26, not a 25/29.
+We then report the mean across DFAs and 95% bootstrap confidence intervals.
 5.1 Sequence Completion
@@
 Of course, BruteForceS is extremely computationally expensive, and, as such, we primarily focus on the $n$-GramS heuristics as our baselines.
-Still, we find, similarly to the transducer task, that $n$-GramS heuristics tend to outperform LLMs.
+Still, we find that $n$-GramS heuristics tend to outperform LLMs.
 As seen in Table 2, we find that giving the model the opportunity to logically reason about the prompt via chain-of-thought and present a conclusion has inconsistent results.
@@
 One potential problem with using this task for cross-model comparisons is the relevance of tokenization.
-Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix B for details.
+Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix D for details.
 Model Basic More-Expl COT Red-Green
@@
 claude-3.5 86.9 (83.3–90.0) 87.1 (83.9–90.2) 76.4 (72.9–79.9) 82.9 (78.9–86.9)
-Table 2: Results for chat models (where we investigated multiple prompts).
-For all other models, we only used the Basic prompt.
-We find that non-COT prompts consistently work better for the Transducer task, with more mixed results on sequence completion.
+Table 2: Results for models where we investigated multiple prompts (we only used Basic on other models).
+We bold the best prompt for each model.
+Non-COT prompts consistently work better for the Transducer task, with more mixed results on sequence completion.
 5.2 Transducer
 Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ model only achieving 96.4% accuracy.
-However, relative results should still be valid as the impossible problems are present with equal probability for all models.
+However, relative results should still be valid as the impossible problem instances are present with equal probability for all models.
 We find that in general all LLMs underperform a 4-GramT model, demonstrating that they are unable to adequately solve this task.
-The relative performance of the models also does not generally correspond to their overall scale, with open source LLama-3 and Mistral Nemo 8B parameter models outperforming Claude and GPT-4o.
-Even within a model class we find no clear pattern: GPT-4o is outperformed by GPT 3.5, Llama 3-70B has similar performance to Llama 3-8B, and the Mistral Nemo 12B models perform similarly to Nemo Minitron 8B.
-Coding models also demonstrate no particular advantage on this task.
+The relative performance of the models also does not correspond to their overall scale, with open source LLama-3 and Mistral Nemo 8B parameter models outperforming Claude and GPT-4o.
+Even within a model class we find no clear pattern: GPT-4o and o1-preview [^4] The particularly poor performance of o1-preview may be due to the model not supporting temperature 0.
+In other experiments, we found that GPT models with temperature 1 tended to perform poorly.
+See Appendix B for more details.
+are outperformed by GPT 3.5, Llama 3-70B has similar performance to Llama 3-8B, and the Mistral Nemo 12B models perform similarly to Nemo Minitron 8B.
+Coding models also demonstrate no advantage on this task.
 The generally lower performance of chat-oriented models suggests this task is better suited to non-chat models.
-To investigate that this is not specific to the Basic prompt, we investigate several chat models on the other prompts.
-As seen in Table 2, our chain-of-though and word problem prompts, which attempt to leverage the full reasoning capabilities of chat models, also fare poorly, performing similarly or worse to the Basic prompt on the Transducer task.
+To investigate that this is not specific to the Basic prompt, we investigate other prompts for chat models.
+As seen in Table 2, our chain-of-though and word problem prompts, which attempt to leverage the full reasoning capabilities of chat models, also fare poorly, performing similarly or worse to the Basic prompt on the Transducer task in all cases.
 Overall, we conclude that LLMs are unable to perform the DFA transducer inference task to a reasonable degree.
@@
 Instead, it seems the LLMs are unable to detect patterns when those patterns are drawn from an unfamiliar source, even a relatively simple one.
+5.3 Comparison of Benchmarks
+Figure 3: Transducer and sequence completion results plotted against each other.
+Points represent the mean over several DFAs and intervals represent 95% confidence intervals.
+Points are colored by model type, with the best and worst model by each metric in each category labeled, as well as all baseline and proprietary models.
+Figure 3 displays the relationship between model performance on the Sequence Completion and Transducer benchmarks.
+While at a high level, there is a positive correlation between the two, there are a few notable differences.
+For one, the Code models perform notably better than other open source models on Sequence Completion, but not on Transducer.
+Additionally, on Transducer, a ceiling on performance is observed, where LLMs cluster together between 3-GramT and 4-GramT performance; this clustering does not appear on the Sequence Completion benchmark.
 5.4 Case Study: Sum Modulo 3 DFA
@@
 1b and 1c lead to 0 14 17
-2 red rooms 2 3
 2-periodic 9 14
 3-periodic 4 4
-$>3$-Gram 4 5
-Sum Modulo 3 0 0
-Figure 3: Results on Sum Modulo 3 DFA.
-(a) Venn diagram results for CR (claude-3.5 using Red-GreenT), QB (qwen-2.5-coder-instruct-7B using BasicT), and 6-GramT.
-Percentages are accuracy numbers, and venn diagram is error counts, out of 1000.
-(b) Results of qualitative analysis.
-30 CR reasoning traces were sampled in each condition, with Incorrect traces guaranteed to correspond to problems 6-GramT got correct.
-We focus on the sum modulo 3 DFA in this section.
-We depict this DFA in Figure 1.
+2 red rooms 2 3
+Figure 4: Results on Sum Modulo 3 DFA. (a) Venn diagram of errors (out of 1000). Labeled percentages are accuracies. (b) Results of qualitative analysis, out of 30 in both cases.
+We investigate the transducer task on the DFA depicted in Figure 1.
 This DFA can be interpreted as an arithmetic check, where a represents 0, b represents 1, and c represents 2, and the DFA accepts strings whose sum is equal to 0 modulo 3.
-We focus on the Transducer setting as sum modulo 3 is nearly impossible (for LLMs and $n$-Gram models alike) in the sequence completion setting due to the issues raised in Section 3.3.
-For this case study, we focus on the model/prompt combinations QB and CR, defined as
-  • QB: qwen-2.5-coder-instruct-7B with BasicT. Selected as it is the best performing combination overall.
-  • CR: claude-3.5 with Red-GreenT. Selected as it is the best performing combination that provides an explanation (needed later for our qualitative analysis)
+For this case study, we focus on the model/prompt combinations MB and CR, defined as
+  • MB: mistral-nemo-minitron-8B/BasicT. Selected as it is the best performing combination overall.
+  • CR: claude-3.5/Red-GreenT. Selected as it is the best performing combination that provides an explanation (needed later for our qualitative analysis)
 Figure 4a depicts the number of errors each model receives on 1000 instances of the transducer task for this DFA.
@@
 […]
-Despite transformers’ theoretically high computational capacity, without the ability to pattern match to existing problems, Claude uses an unsophisticated and ineffectual approach.
+Despite transformers’ high computational capacity, without the ability to pattern match to existing problems, Claude uses an unsophisticated and ineffectual approach.
 6 Conclusion
@@
 While our conclusions are limited by the finite nature of our prompt set, this suggests that they do, in fact, possess some latent understanding of language, but this understanding is inferior to basic n-gram models for $n>3$.
-These results expose a particular weakness in LLMs, as many potential foundation model applications involve tasks that are not expressed in familiar human languages or pre-existing programming languages.
+Many potential foundation model applications involve tasks that are not expressed in familiar human languages or pre-existing programming languages.
 More specifically, in tasks where there is a need to produce an output in a precise, atypical, format, we should be skeptical of the ability of LLMs to in-context-learn this format.
 For these tasks, it may be prudent to seek a new approach.
-Acknowledgments
-We thank Marc Marone and Theo Olausson for their advice on experiment design.
+Impact Statement
+Aside from the social consequences of this work as related to advancing the field of Machine Learning in general, this work has the goal of advancing the field of benchmarks in Machine Learning.
+While we view this as a positive objective, as it ensures that models are being evaluated fairly, it might have negative consequences insofar as benchmarking techniques might be best left unpublished to prevent deliberate or unintentional overfitting.
 Appendix A Details on Sampling
@@
 We also randomly assign each state to be accept or reject with probability 50%.
-We then reject any DFA that has all accept or all reject states, or for which certain states are unreachable from the start state.
+We then reject any DFA that has all accept or all reject states (so only DFAs with 1 or 2 accept states are allowed), or for which certain states are unreachable from the start state.
 A.2 Sampling of Sequence Completion Tasks
@@
 Each example sequence is sampled uniformly from the space of $\{a,b,c\}^{10}$ and then rejected if the DFA does not accept the sequence.
-Our distinct prefix and completion are sampled uniformly from $\{a,b,c\}^{5}\times\{a,b,c\}^{5}$, and is rejected if the DFA does not accept the concatenation of the two, or if the prefix is the prefix of any of the previous sequences.
+Our distinct prefix and completion are sampled uniformly from $\{a,b,c\}^{5}\times\{a,b,c\}^{5}$, and are rejected if the DFA does not accept the concatenation of the two, or if the prefix is the prefix of any of the previous sequences.
 We then discard the completion.
 If we, at any point, reject 50 sequences when attempting to sample a sequence or prefix, we return an error.
-We run a “pilot” sampling for a DFA to ensure that it is valid, in which we sample an instance as described abov.
+We run a “pilot” sampling for a DFA to ensure that it is valid, in which we sample an instance as described above.
 If there is an error in sampling this pilot instance, we reject the DFA.
@@
 At this stage, if there is an error in sampling, we reject the instance rather than the DFA.
+This pilot sample rejection procedure leads to a slight bias towards 2-accept state DFAs over 1-accept state DFAs, as measured by the $\textsc{Random}_{S}$ baseline.
 A.3 Sampling of Transducer Tasks
@@
 If every transducer trace ends with a 0 or every trace ends with a 1, we reject the DFA and resample.
+Appendix B Results of o1-preview
+We evaluated o1-preview on 10 DFAs, using 30 problem instances per DFA of the Transducer task, as in other Transducer experiments, and the $\textsc{Basic}_{T}$ prompt, as this is the most neutral prompt.
+Table 3 displays results on each DFA.
+Overall, while these results are not on a particularly large sample, they fairly definitively demonstrate that o1-preview does not achieve strikingly good performance on this task.
+DFA o1-preview gpt-4o 6-Gram
+1 25/30 27/30 26/30
+2 23/29 24/30 25/30
+3 19/30 23/30 28/30
+4 22/30 23/30 28/30
+5 29/29 30/30 30/30
+6 19/30 24/30 30/30
+7 17/29 23/30 25/30
+8 23/30 25/30 26/30
+9 21/30 28/30 30/30
+10 29/30 29/30 30/30
+Table 3: Results on each DFA.
+We find that in 7 cases, o1-preview underperforms gpt-4o, in 2 cases it gets the same number of instances wrong but provides a non-answer on an additional instance, and in 1 case it ties gpt-4o.
+In no cases does it outperform.
+Appendix C More details on Sum Modulo 3 DFA case study
+Figure 5 depicts the results of the Sum Modulo 3 experiment, but filtered for two conditions.
+In the (a) condition, the trace ends in such a way that a 3-Gram model would be able to determine the output, and the (b) condition is the complement.
+Figure 5: Results on Sum Modulo 3 DFA under trivial / nontrivial conditions.
+Percentages are accuracy numbers, and venn diagram is error counts.
+(a) In this condition, CR and the 6-GramT both get very high accuracies, with nearly all 6-GramT also being CR errors.
+MB does relatively poorly.
+(b) In this condition, models do significantly more poorly overall, with CR in particular performing worse than chance.
+Here, errors are more symmetric, with more 6-GramT errors that are not accounted for by either or both model, indicating that a larger fraction of both successes and failures in this condition are down to random chance.
 Model BasicS Basic-CommasS
@@
 Table 4: Results on Sequence Completion Task. We compare BasicS to the comma-variant Basic-CommasS.
-Appendix B Sequence Completion task prompt with Commas
+Appendix D Sequence Completion task prompt with Commas
 To avoid tokenization differences with models, we also investigate a version of our Sequence Completion prompt that uses spaces and commas between the elements of the sequence.
@@
 All are worse with commas than without.
-Appendix C More details on Sum Modulo 3 DFA case study
-Figure 4: Results on Sum Modulo 3 DFA under trivial / nontrivial conditions.
-Percentages are accuracy numbers, and venn diagram is error counts.
-(a) In this condition, CR and the 6-GramT both get very high accuracies, with nearly all 6-GramT also being CR errors.
-QB does poorly.
-(b) In this condition, models do significantly more poorly overall, with CR in particular performing worse than chance.
-Here, errors are more symmetric, with more 6-GramT errors that are not accounted for by either or both model, indicating that a larger fraction of both successes and failures in this condition are down to random chance.
-Appendix D Prompt Listings
+Appendix E Prompt Listings
 Prompt $T$ $S$
@@
 Table 5: Shortened summary of each prompt
-D.1 Summaries
+E.1 Summaries
 Table 5 contains summaries of each prompt.
-D.2 Full example listings
-D.2.1 BasicT
+E.2 Full example listings
+E.2.1 BasicT
 You are a sequence completion model. Output the next element of the sequence, and nothing else.
 a, 1, b, 1, a, 1, b, 1, b, 1, c, 0, a, 1, c, 1, a, 1, a, 1, a, 1, c, 1, b, 1, c, 0, c, 1, a, 1, b, 1, b, 1, b, 1, b, 1, a, 1, b, 1, a, 1, a, 1, b, 1, c, 0, a, 1, c, 1, a, 1, b,
-D.2.2 More-ExplT
+E.2.2 More-ExplT
 You are a sequence completion model.
@@
 a, 1, b, 1, a, 1, b, 1, b, 1, c, 0, a, 1, c, 1, a, 1, a, 1, a, 1, c, 1, b, 1, c, 0, c, 1, a, 1, b, 1, b, 1, b, 1, b, 1, a, 1, b, 1, a, 1, a, 1, b, 1, c, 0, a, 1, c, 1, a, 1, b,
-D.2.3 COTT
+E.2.3 COTT
 A DFA is a finite-state machine that accepts or rejects a given string of symbols, by running through a n-state sequence uniquely determined by the string.
@@
 Output sequence: 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1,
-D.2.4 Red-GreenT
+E.2.4 Red-GreenT
 You are in a house of rooms and portals.
@@
 Then, you walk through a portal labeled "B" and end up in a ...
-D.2.5 BasicS
+E.2.5 BasicS
 The following strings come from an alien language that follows a simple grammar.
@@
 caacb
-D.2.6 Basic-CommasS
+E.2.6 Basic-CommasS
 The following strings come from an alien language that follows a simple grammar.
@@
 c, a, a, c, b,
-D.2.7 More-ExplS
+E.2.7 More-ExplS
 I have a 3-state DFA model that outputs either 0 or 1 after each element I input.
@@
 caacb
-D.2.8 COTS
+E.2.8 COTS
 I have a 3-state DFA model that outputs either 0 or 1 after each element I input.
@@
 caacb
-D.2.9 Red-GreenS
+E.2.9 Red-GreenS
 You are outside a house of rooms and portals.

### v4

@@
 Can LLMs pick up language structure from examples?
-Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures and vocabularies.
+Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures.
 However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with seen languages.
@@
 1 Introduction
-Figure 1: We use (a) randomly generated deterministic finite automata (DFAs) to sample previously unseen grammars to test models with.
-The DFA shown here, modeling the sum modulo 3 operation (with abc representing 0, 1, and 2 respectively), can be used to accept or reject strings from a 3-character alphabet.
+Figure 1: We sample randomly generated languages to test LLMs by sampling deterministic finite automata (DFAs).
+(a) The DFA shown here, modeling the sum modulo 3 operation (with abc representing 0, 1, and 2 respectively), can be used to accept or reject strings from a 3-character alphabet.
 Accepted strings belong to the grammar, and rejected strings do not.
@@
 Therefore, ensuring that even an entirely novel invented problem is not a variation on a theme is intractable.
-To circumvent this problem, we propose the following general approach: first define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level.
-Then, sample random languages from this space.
-By sampling randomly, we can guarantee no bias towards canonical languages that might share structure with ones in the training dataset.
+To circumvent this problem, we propose the following general approach: first we define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level.
+Then, we sample random languages from this space.
+By sampling randomly, we can guarantee no bias towards canonical languages that might share structure with common ones in the training dataset.
 In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level, but this technique can be generalized to produce benchmarks of any difficulty level.
@@
 Yao & Koller (2022) show that smaller language models like BART and T5 can struggle on these tasks, but to our knowledge there are not comprehensive experiments extending this line of work to LLMs.
+2.4 Training transformers on Formal Languages
+Vafa et al.
+(2025) frame world modeling as a latent DFA identification task, finding that transformers trained on DFA traces (of massive DFAs representing board games and city maps) do not reconstruct the underlying DFA.
+Akyürek et al.
+(2024) find that transformers trained on 4-12 state DFA transducer traces more effectively learn to in-context-learn regular languages than RNNs.
+Here, we evaluate much larger and more capable LLMs on much simpler 3-state DFAs.
+However, our goal is evaluating the abilities of foundation models (trained on a broad corpus rather than a tailored one), and is thus not strictly easier.
 3 DFA Reasoning Tasks
@@
 Examples of languages recognized by DFAs include simple ones like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7.
-Finite automaton languages are equivalent to regular languages, which are recognized by regular expressions.
-One simple metric of the difficulty of a regular language is the number of states in the corresponding DFA, which represents the amount of memory the automaton has at any point while processing a given sequence.
+Type 3 languages are also known as regular languages, which are recognized by regular expressions.
+One simple metric of the difficulty of a regular language is the number of states in the corresponding DFA, i.e., the amount of working memory.
 [^1] There are other metrics of difficulty, but we choose number of states as it is highly parsimonious.
-DFAs with 2 states have the property that their set of states is no larger than the output set $\{0,1\}$, and, therefore, do not have any hidden state.
+2-state DFAs have the property that their set of states is no larger than the output set $\{0,1\}$, and, therefore, do not have any hidden state.
 We thus explore 3-state DFAs, as this is the simplest nontrivial case.
@@
 We first pose a sequence completion task, in which models must complete a sequence in a given DFA’s language.
-We study this task primarily because, in practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
+In practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
 To generate test cases for this task given a DFA, we first (1) sample 30 example sequences of length 10 that this DFA accepts, and then (2) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept.
@@
 For details on sampling, see Appendix A.2.
-We evaluate models by (1) sampling a DFA, (2) sampling 30 problem instances at random (each of which contains 30 example sequences and a distinct prefix), and then (3) computing a binary prediction score (whether or not the predicted completion creates a valid string in the language) for each instance separately.
+We evaluate models by (1) sampling a DFA, (2) sampling 30 problem instances at random (each of which contains 30 example sequences and a distinct prefix), and then (3) computing a binary prediction score (whether or not the predicted completion creates a valid string in the language) for each instance separately, then computing a correctness metric as a fraction.
 We then average this metric over several sampled DFAs to produce our accuracy score.
@@
 In this task, an input sequence is annotated with an output at each token, the final output is masked, and the masked output is predicted by a language model.
-We call this a transducer task, as the DFA is used as a machine that converts a sequence of inputs into a sequence of outputs.
+We call this a transducer task, as the DFA converts a sequence of inputs into a sequence of outputs.
 E.g., given the language does the string have an even number of ’a’ tokens and the input abcabcaabbccaa, the annotated string (all that is provided to the model) is a0b0c0a1b1c1a0a1b1b1c1c1a0a and the output to predict is 1.
 For each problem instance, we provide 30 symbols, and for the first 29, the corresponding transducer output.
-This task is significantly more transparent than the sequence completion task as the model has access to intermediate transducer outputs, an (imperfect) proxy for intermediate state. n
+This task is significantly more transparent than the sequence completion task as the model has access to intermediate outputs, an (imperfect) proxy for intermediate state. n
 3.4 Baselines
@@
 We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts.
-Full listings of these prompts can be found in Appendix E, with Table 5 containing a summary of each prompt.
+Full listings of these prompts can be found in Appendix F.
 While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe our set of prompts captures common prompting strategies.
@@
 One potential problem with using this task for cross-model comparisons is the relevance of tokenization.
-Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix D for details.
+Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix E for details.
 Model Basic More-Expl COT Red-Green
@@
 5.2 Transducer
-Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ model only achieving 96.4% accuracy.
-However, relative results should still be valid as the impossible problem instances are present with equal probability for all models.
+Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ baseline only achieving 96.4% accuracy.
+Comparisons are still be valid as all models see the same fraction of unsolvable instances.
 We find that in general all LLMs underperform a 4-GramT model, demonstrating that they are unable to adequately solve this task.
@@
 As seen in Table 2, our chain-of-though and word problem prompts, which attempt to leverage the full reasoning capabilities of chat models, also fare poorly, performing similarly or worse to the Basic prompt on the Transducer task in all cases.
-Overall, we conclude that LLMs are unable to perform the DFA transducer inference task to a reasonable degree.
-This failure cannot be attributed to a lack of world modeling ability, as $n$-GramT models do not construct world models.
+We conclude that LLMs are unable to perform well on the DFA transducer inference task.
+This failure cannot be attributed to a lack of world modeling, as $n$-GramT models do not construct world models.
 Instead, it seems the LLMs are unable to detect patterns when those patterns are drawn from an unfamiliar source, even a relatively simple one.
@@
 Correct Incorrect
-Total 30 30
-a is no-op 21 22
-1b and 1c lead to 0 14 17
-2-periodic 9 14
-3-periodic 4 4
-2 red rooms 2 3
-Figure 4: Results on Sum Modulo 3 DFA. (a) Venn diagram of errors (out of 1000). Labeled percentages are accuracies. (b) Results of qualitative analysis, out of 30 in both cases.
+Total 100% 100%
+a is no-op 70% 73%
+1b and 1c lead to 0 47% 57%
+2-periodic 30% 47%
+3-periodic 13% 13%
+2 red rooms 7% 10%
+Figure 4: Results on Sum Modulo 3 DFA.
+(a) MB=mistral-nemo-minitron-8B/BasicT, CR=claude-3.5/Red-GreenT.
+Venn diagram of errors (out of 1000).
+Labeled percentages are accuracies.
+(b) Results of qualitative analysis, out of 30 in both cases.
 We investigate the transducer task on the DFA depicted in Figure 1.
@@
 Figure 4a depicts the number of errors each model receives on 1000 instances of the transducer task for this DFA.
-In general, nearly all errors made by the 6-GramT model were also made by at least one LLM model, while the two LLM models often made unique errors.
-Interestingly, while this task is better-known than most DFAs, we find that all 3 models perform worse on this DFA than their average across random DFAs.
+Nearly all errors made by the 6-GramT model were also made by at least one LLM, while the two LLMs often made unique errors.
+While this task is better-known than most DFAs, all 3 models perform worse on this DFA than average.
 We also performed a qualitative analysis, investigating CR’s outputs on the Red-GreenT prompt to see what kind of reasoning it is using; specifically we sampled 30 examples where it had the correct answer, and 30 examples where it had the incorrect answer but the 6-GramT model had the correct answer.
@@
 Specifically, it is able to learn that a does not change the output, and that b and c will lead a 1 state to a 0 state.
-These results comport with the overall finding of Table 1, where we found that 3-GramT was the largest $n$-GramT that any LLM outperformed.
+These results comport with the overall finding of Table 1, where we found that 3-GramT was the largest $n$-GramT that any LLM outperformed, as well as our finding that LLM performance decreases for tasks that are not solvable by $n$-Grams; see Appendix C for details.
 […]
@@
 In no cases does it outperform.
-Appendix C More details on Sum Modulo 3 DFA case study
+Appendix C Transducer results by difficulty class
+Figure 5: Transducer results by difficulty class.
+We classify each DFA based on which of the baselines first achieves a score of 28/30 on the given instances.
+6-Gram is excluded as it has very similar performance to 5-Gram.
+Each model’s best prompt results are plotted, with 95% confidence intervals, for all models with at least 100 DFAs; those with 10 or 30 had error bars too large to make this analysis useful.
+Figure 5 displays results by difficulty level, as judged by the smallest $n$-Gram model that can solve a particular task.
+All models behave roughly monotonically, performing more poorly as difficulty increases.
+Additionally, we find that the best models continue to perform similarly to 4-Gram for tasks that 4-Gram does not perfectly solve.
+Appendix D More details on Sum Modulo 3 DFA case study
 Figure 6 depicts the results of the Sum Modulo 3 experiment, but filtered for two conditions.
@@
 Table 4: Results on Sequence Completion Task. We compare BasicS to the comma-variant Basic-CommasS.
-Appendix D Sequence Completion task prompt with Commas
+Appendix E Sequence Completion task prompt with Commas
 To avoid tokenization differences with models, we also investigate a version of our Sequence Completion prompt that uses spaces and commas between the elements of the sequence.
@@
 All are worse with commas than without.
-Appendix E Prompt Listings
+Appendix F Prompt Listings
 Prompt $T$ $S$
@@
 Table 5: Shortened summary of each prompt
-E.1 Summaries
+F.1 Summaries
 Table 5 contains summaries of each prompt.
-E.2 Full example listings
-E.2.1 BasicT
+F.2 Full example listings
+F.2.1 BasicT
 You are a sequence completion model. Output the next element of the sequence, and nothing else.
 a, 1, b, 1, a, 1, b, 1, b, 1, c, 0, a, 1, c, 1, a, 1, a, 1, a, 1, c, 1, b, 1, c, 0, c, 1, a, 1, b, 1, b, 1, b, 1, b, 1, a, 1, b, 1, a, 1, a, 1, b, 1, c, 0, a, 1, c, 1, a, 1, b,
-E.2.2 More-ExplT
+F.2.2 More-ExplT
 You are a sequence completion model.
@@
 a, 1, b, 1, a, 1, b, 1, b, 1, c, 0, a, 1, c, 1, a, 1, a, 1, a, 1, c, 1, b, 1, c, 0, c, 1, a, 1, b, 1, b, 1, b, 1, b, 1, a, 1, b, 1, a, 1, a, 1, b, 1, c, 0, a, 1, c, 1, a, 1, b,
-E.2.3 COTT
+F.2.3 COTT
 A DFA is a finite-state machine that accepts or rejects a given string of symbols, by running through a n-state sequence uniquely determined by the string.
@@
 Output sequence: 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1,
-E.2.4 Red-GreenT
+F.2.4 Red-GreenT
 You are in a house of rooms and portals.
@@
 Then, you walk through a portal labeled "B" and end up in a ...
-E.2.5 BasicS
+F.2.5 BasicS
 The following strings come from an alien language that follows a simple grammar.
@@
 caacb
-E.2.6 Basic-CommasS
+F.2.6 Basic-CommasS
 The following strings come from an alien language that follows a simple grammar.
@@
 c, a, a, c, b,
-E.2.7 More-ExplS
+F.2.7 More-ExplS
 I have a 3-state DFA model that outputs either 0 or 1 after each element I input.
@@
 caacb
-E.2.8 COTS
+F.2.8 COTS
 I have a 3-state DFA model that outputs either 0 or 1 after each element I input.
@@
 caacb
-E.2.9 Red-GreenS
+F.2.9 Red-GreenS
 You are outside a house of rooms and portals.

### v6

@@
-Randomly Sampled Language Reasoning Problems Reveal Limits of LLMs
+Randomly Sampled Language Reasoning Problems Elucidate Limitations of In-Context Learning
 Abstract
-Can LLMs pick up language structure from examples?
-Evidence in prior work seems to indicate yes, as pretrained models repeatedly demonstrate the ability to adapt to new language structures.
-However, this line of research typically considers languages that are present within common pretraining datasets, or otherwise share notable similarities with seen languages.
-In contrast, in this work we attempt to measure models’ language understanding capacity while circumventing the risk of dataset recall.
-We parameterize large families of language tasks recognized by deterministic finite automata (DFAs), and can thus sample novel language reasoning problems to fairly evaluate LLMs regardless of training data.
-We find that, even in the strikingly simple setting of 3-state DFAs, LLMs underperform unparameterized $n$-Gram models on both language recognition and synthesis tasks.
-These results suggest that LLMs struggle to match the ability of basic language models in recognizing and reasoning over languages that are sufficiently distinct from the ones seen at training time, underscoring the distinction between learning individual languages and possessing a general theory of language.
-Machine Learning, ICML
+While LLMs have revolutionized the field of machine learning due to their high performance on a strikingly wide range of problems, they are also known to hallucinate false answers and underperform on less canonical versions of the same tasks.
+There are several emerging theories of LLM performance, among them that LLMs lack world modeling ability, that they have an undesirable bias towards an autoregressive prior, and that they struggle on more novel problems.
+The existing literature on LLM input novelty has focused on tasks of relatively high complexity, studying perturbations of canonical but complex problems.
+In this paper, we attempt to minimize complexity in order to isolate novelty as a factor in LLM underperformance and investigate the power of in-context-learning.
+To this end, we consider an extremely simple domain: next token prediction on simple language tasks.
+The twist is that these language tasks are wholly unseen, as they are randomly drawn from a large, parsimoniously defined set of languages arising from simple grammar rules.
+This experimental setup allows us to evaluate ICL independently of models’ parametric knowledge.
+We find that LLMs uniformly underperform n-gram models on this task, both when used as next token predictors and in chain-of-thought.
 1 Introduction
@@
 (a) The DFA shown here, modeling the sum modulo 3 operation (with abc representing 0, 1, and 2 respectively), can be used to accept or reject strings from a 3-character alphabet.
-Accepted strings belong to the grammar, and rejected strings do not.
-We evaluate models on their ability to (b) act as a transducer, recognizing strings that belong to the DFA-defined grammar, and (c) generate new strings following the grammar.
-Contemporary LLMs have proven themselves to be highly sophisticated natural language completion models that demonstrate many properties of reasoning engines.
-This has prompted questions surrounding the true intelligence of these models, with some arguing that they possess inherent language learning capabilities (Millière, 2024).
-In this paper, we explore the question of whether LLMs have the reasoning capacity to understand the structure of a new language.
-Specifically, we are interested in problems where a model is given a small set of examples from a language and either generates a new sample or determines whether a new sample is from the language or not.
-Some work suggests that LLMs broadly understand language structure because they are able to produce syntactically correct samples from languages they have not been trained on (Athiwaratkun et al., 2022), although LLMs’ performance on low-resource languages tends to be lower than their ability on higher resource languages (Bogin et al., 2023).
-On the other hand, some critics of LLMs argue that LLMs cannot possess an understanding of language structure as they have learned from data rather than possessing a priori universal grammar (Chomsky et al., 2023).
-We do not view learning from data as a fundamental limitation, but we are concerned with the possibility that a language model might only be able to understand linguistic structures similar to those it has seen in training data.
-To distinguish between these possibilities, we wish to evaluate LLMs on a benchmark that considers wholly novel languages, eliminating the possibility of dataset leakage.
-Ideally, models would be tested on a set of language reasoning problems disjoint from data seen during training and validation.
-However, as training datasets for LLMs are generally closed and incredibly vast, human-generated problems in natural language are likely to at least partially overlap in syntax or concept with content LLMs have already seen, making probing a model for its ability to reason about the structure of natural languages nearly impossible.
-Additionally, the problem of determining whether two tasks are semantically identical is itself a nontrivial one.
-Therefore, ensuring that even an entirely novel invented problem is not a variation on a theme is intractable.
-To circumvent this problem, we propose the following general approach: first we define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level.
+Accepted strings belong to the grammar; rejected strings do not.
+We evaluate models on their ability to (b) act as a transducer, recognizing strings that belong to the grammar, and (c) generate new strings following the grammar.
+One of the surprising capabilities of contemporary LLMs is their ability to perform in-context-learning (ICL), in which they learn by demonstration from provided input/output examplars to produce an appropriate output for a new but similarly structured input.
+This capability allows LLMs to generalize to tasks beyond interpolations of their training corpus, resulting in remarkably strong generalization capabilities.
+It is challenging to isolate the effects of ICL, as, given the large datasets LLMs are trained on, it is difficult to distinguish a model detecting novel patterns in examples from it being guided towards knowledge it gained during training.
+Akyürek et al.
+(2022) demonstrate that, in specially trained transformers, ICL is “true learning”.
+However, it remains unclear to what degree the performance of foundation-model LLMs can be attributed to strong ICL capabilities.
+The task best suited to isolating LLM ICL performance should have the following properties: (1) It should be a language completion task within the expressive power of an LLM.
+LLMs are capable of many tasks, but they are primarily models of language, and as such, language tasks are the most fair evaluations.
+(2) It should not require sophisticated world modeling to solve.
+This helps us eliminate a possible source of underperformance distinct from ICL capability.
+(3) It should be selected at random in an unbiased manner to reduce the effect of bias from the training corpus.
+To satisfy these properties, we propose the following general approach.
+First we define a large, exhaustive, and parsimoniously-defined space of languages that represents all languages of a certain difficulty level.
 Then, we sample random languages from this space.
 By sampling randomly, we can guarantee no bias towards canonical languages that might share structure with common ones in the training dataset.
-In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level, but this technique can be generalized to produce benchmarks of any difficulty level.
-Experimental results using this approach suggest that contemporary LLMs possess less sophisticated language pattern recognition abilities than expected; underperforming basic, parameter-free n-gram language models on even the simplest languages.
-These results, combined with LLMs’ impressive results on a variety of specific tasks, suggest that LLMs function as ensemble models over language tasks they have seen in their dataset, but do not possess the ability to generalize to entirely novel language reasoning tasks.
+In this work, we use languages recognized by 3-state DFAs as these are the lowest nontrivial difficulty level.
+[^1] This technique can be generalized to produce benchmarks of any difficulty level.
+For larger numbers of states, we would be able to guarantee that the majority of the exponentially large number of corresponding languages do not lie in the training dataset by a pigeonhole argument; unfortunately this does not apply to the relatively small set of 3-state DFAs (there are only 78786).
+However, they still represent an set of tasks of a particular difficulty level not biased towards the canonical.
+Finally, to ensure we are not measuring world modeling performance, we compare to $n$-Gram baselines that are not capable of anything other than matching clusters of tokens.
+Our results demonstrate that even for very simple language induction tasks that don’t rely on world modeling or background knowledge, LLM ICL still underperforms simple language models when dealing with randomly sampled and likely unfamiliar problem instances.
+These results suggest that while LLMs can pick up some learning signal from examples in prompts, this in-context learning is not competitive with even very primitive forms of learning, and suggests that LLMs do not posses the ability to generalize to entirely novel language reasoning tasks.
 In summary, we make the following contributions:
   1. 1.
-    We introduce a benchmark for LLM language reasoning evaluation, disjoint from natural language web data.
+    We introduce a benchmark for LLM ICL language reasoning evaluation, consisting of novel tasks.
   2. 2.
@@
 2 Related Work
+LLMs are known to fail in many cases, with some suggesting that these failures are due to lack of a world model (Valmeekam et al., 2022) or “embers” of autoregression polluting non-autoregressive task performance (McCoy et al., 2023).
+Another theory is that of task novelty; that is, LLMs perform worse on tasks more dissimilar from those seen during training.
+2.1 Language Understanding and LLMs
+LLMs can be quite adept at generating programs in general-purpose programming languages (Xu et al., 2022a).
+In contrast, adapting models to understand domain-specific languages (Mernik et al., 2005) introduces unique problems such as navigating idiosyncratic syntax and semantics and leveraging sparse sample language data.
+To address these challenges, researchers have considered how well general-purpose LLMs can use language reasoning skills to quickly understand rare or unseen DSLs with only a small set of exemplars (Joel et al., 2024).
+While most work in this vein focuses on semantic parsing for downstream applications (Lin et al., 2023), selecting exemplars (Zhao et al., 2021), and improving DSL recognition by leveraging more common languages (Bogin et al., 2023), experiments show strong baseline performance for LLM DSL recognition and parsing out-of-the-box (Wang et al., 2024).
+Some have suggested that indicate that LLMs may possess emergent language reasoning abilities (Millière, 2024).
+Related lines of work are compositional generalization (Xu et al., 2022b), which assesses models’ ability to organize known units into novel structures, and structural generalization (Yao & Koller, 2022), which assesses models’ ability to recognize new structures.
+Yao & Koller (2022) show that smaller language models like BART and T5 can struggle on these tasks, but to our knowledge there are not comprehensive experiments extending this line of work to LLMs.
 2.2 Reasoning with LLMs
@@
 Patel et al.
-(2021) demonstrate that much of LLM mathematical reasoning can be explained by shallow heuristics, and Razeghi et al.
-(2022) similarly find that term frequency in training data impacts models’ in-context learning ability.
-In comparison to these, we explore the distinction described by Patel et al.
-(2021), but push both language simplicity and language unfamiliarity to their limits, by exploring simple languages recognized by randomly sampled DFAs.
-This enables us to evaluate the ability of LLMs to reason about language.
-2.3 Language Understanding and LLMs
-LLMs can be quite adept at generating programs in general-purpose programming languages (Xu et al., 2022a).
-In contrast, adapting models to understand domain-specific languages (Mernik et al., 2005) introduces unique problems such as navigating idiosyncratic syntax and semantics, and leveraging sparse collections of sample language data.
-To address these challenges, researchers have considered how well general-purpose LLMs can use language reasoning skills to quickly understand rare or unseen DSLs with only a small set of exemplars (Joel et al., 2024).
-While most work in this vein focuses on semantic parsing for downstream applications (Lin et al., 2023), selecting exemplars (Zhao et al., 2021), and improving DSL recognition by leveraging more common languages (Bogin et al., 2023), experiments show strong baseline performance for LLM DSL recognition and parsing out-of-the-box (Wang et al., 2024), indicating that LLMs may possess emergent language reasoning abilities.
-Related lines of work are compositional generalization (Xu et al., 2022b), which assesses models’ ability to organize known units into novel structures, and structural generalization (Yao & Koller, 2022), which assesses models’ ability to recognize new structures.
-Yao & Koller (2022) show that smaller language models like BART and T5 can struggle on these tasks, but to our knowledge there are not comprehensive experiments extending this line of work to LLMs.
+(2021) demonstrate that much of LLM mathematical reasoning can be explained by shallow heuristics, Razeghi et al.
+(2022) similarly find that term frequency in training data impacts models’ in-context learning ability, and (McCoy et al., 2023) theorizes that “embers” of autoregression are polluting non-autoregressive task performance.
+The effect of novelty on performance has also been explored in prior work, generally via investigating perturbations of existing tasks, taking existing problems (that are often inherently quite complex, and are only easy because they are well known, e.g., addition of numbers or logical reasoning over natural language) and changing one small aspect of the problem (Wu et al., 2024; Saparov et al., 2023).
+We also study task novelty but push both language simplicity and language unfamiliarity to their limits, by exploring simple languages recognized by randomly sampled DFAs.
+This enables us to best isolate the power of ICL in the language domain, where LLMs should perform best.
 2.4 Training transformers on Formal Languages
+A key assumption behind this work is that the tasks we are using to evaluate LLMs are solvable by LLMs.
 Vafa et al.
-(2025) frame world modeling as a latent DFA identification task, finding that transformers trained on DFA traces (of massive DFAs representing board games and city maps) do not reconstruct the underlying DFA.
+(2025) frame world modeling (a statistical model inferring the true underlying causal graph behind the data being observed) as a latent DFA identification task, finding that transformers trained on DFA traces (of massive DFAs representing board games and city maps) do not reconstruct the underlying DFA.
+Other work also trains language models on formal languages (Butoi et al., 2024; Bhattamishra et al., 2023; Valvoda et al., 2022) and probabilistic formal languages (Borenstein et al., 2024).
 Akyürek et al.
-(2024) find that transformers trained on 4-12 state DFA transducer traces more effectively learn to in-context-learn regular languages than RNNs.
-Here, we evaluate much larger and more capable LLMs on much simpler 3-state DFAs.
-However, our goal is evaluating the abilities of foundation models (trained on a broad corpus rather than a tailored one), and is thus not strictly easier.
+(2024) find that transformers trained on 4-12 state DFA transducer traces more effectively learn to in-context-learn regular languages than RNNs or $n$-Gram models.
+Therefore, in this work, where we evaluate much larger LLMs on much simpler 3-state DFAs, we can be confident that underperformance relative to $n$-Grams is not linked to inherent transformer limitations and must be instead related somehow to specific properties of foundation models.
 3 DFA Reasoning Tasks
@@
 The original Chomsky Hierarchy (Chomsky, 1959) separates language into four types (Figure 2).
-We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automaton (DFA).
+We focus on the task of understanding Type 3 languages, the simplest form of language in the hierarchy, that are recognized by a Deterministic Finite Automata (DFAs) whose outputs are boolean ({0, 1}).
 Examples of languages recognized by DFAs include simple ones like binary strings with an even number of ones, and even such examples as numbers in base 10 divisible by 7.
@@
 In this work, we focus on the simplest language type in the hierarchy, regular grammars, which are recognized by deterministic finite automata (DFAs).
-3.2 Sequence Completion Task
+3.2 Language Reasoning Tasks
+We define a language reasoning task as a task corresponding to some latent language $\mathcal{L}$, where a set of positive/negative examples is provided to a model, with the goal being either classification of a new string in this language, or completion of an existing string to place it in the language.
+3.2.1 Sequence Completion Task
 We first pose a sequence completion task, in which models must complete a sequence in a given DFA’s language.
-In practice, most language data that models encounter will be in roughly this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed via next token prediction.
-To generate test cases for this task given a DFA, we first (1) sample 30 example sequences of length 10 that this DFA accepts, and then (2) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept.
+This mirrors how foundation models are trained using masked language modeling, where data is presented in this format, with several example sequences in a given language followed by a distinct prefix that needs to be completed.
+To generate test cases for this task given a DFA, we (1) sample 30 example sequences of length 10 that this DFA accepts, and then (2) sample a distinct prefix of length 5 that is not a prefix of any of our 30 example sequences, with the property that there exists some length-$\leq 5$ completion of this prefix that the DFA would accept.
 The task is to find a completion (not necessarily the same completion found in sampling) of this prefix of between 1 and 5 characters such that the DFA accepts the full sequence.
@@
 We then average this metric over several sampled DFAs to produce our accuracy score.
-3.3 Transducer Task
+3.2.2 Transducer Task
 While the sequence completion task is the natural one that comes to mind as a basic language task, it has a difficulty-gap problem.
-Specifically, the issue is that many DFAs, including the one shown in Figure 1, recognize languages that are particularly difficult to identify based on a set of examples, unless you build some kind of world model.
-Other DFAs end up being trivial to generate a completion for by analyzing common suffixes.222 The difficulty gap exists because a set of recognized sequences of length 10 gives no direct insight into intermediate states between the first and tenth token.
+Specifically, many DFAs, including the one shown in Figure 1, recognize languages that are particularly difficult to identify based on a set of examples, unless you build some kind of world model.333 The difficulty gap exists because a set of recognized sequences of length 10 gives no direct insight into intermediate states between the first and tenth token.
 As such, to be able to utilize this information for languages like the one in Figure 1 where there are no “resets” (sequences of symbols that necessarily lead to a particular state), a model must be capable of hollistically evaluating the entire sequence, probably requiring a world model.
 Many other DFAs contain these resets, but do so in such a way that makes it possible to e.g., recognize that all sequences that end in a are in the language, making the problem trivial.
-To provide a more direct evaluation of non-world-modeling-based pattern recognition, we explore the Transducer task.
+This is problematic as we would like to be able to assess the performance of language models at pattern recognition, independent of their world modeling abilities.
+To assess pattern recognition, we explore the Transducer task.
 In this task, an input sequence is annotated with an output at each token, the final output is masked, and the masked output is predicted by a language model.
-We call this a transducer task, as the DFA converts a sequence of inputs into a sequence of outputs.
-E.g., given the language does the string have an even number of ’a’ tokens and the input abcabcaabbccaa, the annotated string (all that is provided to the model) is a0b0c0a1b1c1a0a1b1b1c1c1a0a and the output to predict is 1.
+E.g., given the language even number of ’a’ tokens and the input abcabcaabbccaa, the annotated string (all that is provided to the model) is a0b0c0a1b1c1a0a1b1b1c1c1a0a and the output to predict is 1.
 For each problem instance, we provide 30 symbols, and for the first 29, the corresponding transducer output.
-This task is significantly more transparent than the sequence completion task as the model has access to intermediate outputs, an (imperfect) proxy for intermediate state. n
+This task is significantly more transparent than the sequence completion task as the model has access to intermediate outputs, an (imperfect) proxy for intermediate state.
 3.3 Baselines
-l To contextualize LLM accuracies, waseline models with varying degrees of sophistication.
+To contextualize LLM accuracies, we provide several baselines with varying degrees of sophistication.
 Sequence Completion Task
-For the Sequence Completion task, we have three kinds of baseline.
+For the Sequence Completion task, we have four baselines.
   • RandomS baseline: produce a random string of length 5 characters.
@@
 This baseline measures that bias.
-  • Common-SuffixS baseline: find the completion $s$ of length between 1 and 5 that maximizes ((((# of occurrences as a ${}\times|s|)$.
-This baseline does not take the distinct prefix into account, and instead tries to find a universal completion that will always end in an accept state for this language.
-  • $n$-GramS baseline: we take the last $n-1$ characters of the distinct prefix and search to see if they appear in any of the example sequences at a position where the sequence following is an appropriate length to be a completion (at least 1 but at most 5).
+[…]
+  • $n$-GramS baseline: we take the last n1n-1 characters of the distinct prefix and search to see if they appear in any of the example sequences at a position where the sequence following is an appropriate length to be a completion (at least 1 but at most 5).
 We then take a plurality vote among the completions and return this, breaking ties arbitrarily.
-If there are no matches, we return the result of $(n-1)$-GramS.
-Technically these cover more than $n$ characters, since the completion is often $>1$ character long; for simplicity, however, we keep the naming consistent with the Transducer baselines.
+If there are no matches, we return the result of (n1)(n-1)-GramS.
+Technically these cover more than $n$ characters, since the completion is often >1>1 character long; for simplicity, however, we keep the naming consistent with the Transducer baselines.
+Despite the similarity between an $n$-Gram and a DFA in terms of token-to-token transitions, $n$-Grams do not have access to DFA hidden state and thus cannot solve arbitrary DFA language problems, regardless of $n$.
   • Brute-ForceS: take all possible DFAs with 3 states and 3 symbols.
@@
 Note that these baselines are entirely unparameterized and operate identically regardless of the underlying DFA.
-This makes them direct comparisons to using LLMs in in-context-learning.
+This makes them direct comparisons to using LLMs in in-context-learning [^4] One thing to note is that LLMs are required to determine that they are performing next token prediction on a particular string from a natural language description such as “You are a sequence completion model…,” while $n$-Gram models are programmed to do so.
+However, we believe all LLMs we evaluate are sophisticated enough to accomplish this without issue..
 We do not consider BruteForceS to be a reasonable comparison due to its computational complexity, and instead consider it an upper bound on performance on this particular task.
+We choose $n$-Gram baselines as they are are unambiguously representable by transformers (Svete & Cotterell, 2024), so a transformer model should be able to match their performance.
 Transducer Task
@@
   • NullT baseline: for a given DFA, whichever of the following strategies produces a higher accuracy: always predict 0 or always predict 1.
-  • $n$-GramT baseline: take the $n-1$ symbols ending at the end of the concatenated transducer sequence (e.g., for $n=5$ and the above example, this would be 1a0a).
-If that sequence does not appear elsewhere in the sequence, return the result of the $(n-1)$-GramT baseline.
+  • $n$-GramT baseline: take the n1n-1 symbols ending at the end of the concatenated transducer sequence (e.g., for $n=5$ and the above example, this would be 1a0a).
+If that sequence does not appear elsewhere in the sequence, return the result of the (n1)(n-1)-GramT baseline.
 Otherwise, take the token that appears immediately after each occurrence.
@@
 4 Experiments
-We evaluated the open-source models Llama 3-8B, Llama 3-70B (AI@Meta, 2023), and Llama 3-8B-Instruct (AI@Meta, 2024), Mistral Nemo Minitron 8B (NVIDIA, 2024), Mistral Nemo Base 2407 (Mistral AI, 2024b) and Mistral Nemo Instruct 2407 (Mistral AI, 2024c), Gemma 7B (Google, 2024), and Falcon 7B (Almazrouei et al., 2023).
-We also evaluated the open-source code models StarCoder2-15B (Lozhkov et al., 2024), Codestral-22B-v0.1 (Mistral AI, 2024a), Deepseek Coder 33B Instruct (Deepseek, 2024), Qwen2.5-Coder-7B, Qwen2.5-Coder-7B-Instruct, and Qwen2.5-Coder-32B-Instruct (Hui et al., 2024).
-Finally, we evaluated the proprietary models GPT-3.5-turbo-instruct, GPT-3.5 Chat (turbo-0125) (OpenAI, 2024a), GPT-4o-mini (2024-07-18), GPT 4o (2024-05-13) (OpenAI, 2024b), o1-preview (2024-09-12) (OpenAI et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024).
-For each open source model, we used a local VLLM (Kwon et al., 2023) server for evaluation and always evaluated on 1000 distinct DFAs.
-For GPT-4o and Claude, we evaluated on 30 DFAs due to computation costs.
-For o1-preview we evaluated on only 10 DFAs, and only on the Transducer task (which we felt was a better fit for a reasoning model).
-For gpt-3.5 and gpt-4o-mini, we evaluated on 100 DFAs.
-All models were evaluated with temperature 0, except o1-preview333o1-preview does not allow setting a non-default temperature..
-For both tasks, we consider four prompting formats:
+We evaluated the open-weight models Llama 3-8B, Llama 3-70B (AI@Meta, 2023), Llama 3.1-8B (AI@Meta, 2024b), Llama 3.1-8B-Instruct (AI@Meta, 2024c), Llama 3.1-70B (AI@Meta, 2024a), Mistral Nemo Minitron 8B (NVIDIA, 2024), Mistral Nemo Base 2407 (Mistral AI, 2024b) and Mistral Nemo Instruct 2407 (Mistral AI, 2024c), Gemma 7B (Google, 2024), Falcon 7B (Almazrouei et al., 2023), Qwen 2.5-7B and Qwen 2.5-32B Team (2024).
+We also evaluated the open-weight code models StarCoder2-15B (Lozhkov et al., 2024), Codestral-22B-v0.1 (Mistral AI, 2024a), Deepseek Coder 33B Instruct (Deepseek, 2024), Qwen2.5-Coder-7B, Qwen2.5-Coder-7B-Instruct, and Qwen2.5-Coder-32B-Instruct (Hui et al., 2024).
+Finally, we evaluated the proprietary models Claude 3.5 Sonnet (Anthropic, 2024), GPT-3.5-turbo-instruct, GPT-3.5 Chat (turbo-0125) (OpenAI, 2024a), GPT-4o-mini (2024-07-18), GPT 4o (2024-05-13) (OpenAI, 2024b), o3-mini (2025-01-31) (OpenAI, 2025b), and gpt-5 (2025-08-07) (OpenAI, 2025a).
+For both tasks, we consider two main prompting formats.
 Basic provides no context, presenting the problem as a generic sequence generation or next-token prediction task, where output is provided immediately following the input, with no space to think.
-  • More-Expl explains that the strings are generated from a DFA with 3 states, but is otherwise identical to Basic. This remains a sequence generation/next token prediction task.
-  • COT provides the same information as More-Expl and additionally invokes chain-of-thought reasoning to help the model reason over the task.
-Here, the model is given space to reason before providing a tagged answer.
+Basic-COT provides the same prompt but asks the model to think step by step and provide an answer.
+These prompts test ICL, presenting the task in an unstructured manner and requiring the model to learn the problem structure via induction.
+Our main results are the maximum over these two prompting strategies.
+We also provide three “control” prompting formats where information on the problem structure is provided.
+More-Expl explains that the strings are generated from a simple grammar, but is otherwise identical to Basic.
+This remains a sequence generation/next token prediction task.
+DFA-COT provides the full structure of the latent language, stating that it is a 3-state DFA, and additionally invokes chain-of-thought reasoning to help the model reason over the task.
 Red-Green casts the tasks as independent word problems that describe the underlying grammar structure without relying on world knowledge about DFAs and regular languages.
 It describes an N-state DFA as a house with N rooms, each of which has 3 portals that deterministically go to other rooms (or back to the same room), where the walls of each room are red or green (mirroring transducer output symbols 0 and 1).
-Similarly to COT, the model is given space to show work before providing a tagged answer.
+Similarly to DFA-COT, the model is given space to show work before providing a tagged answer.
 We produce versions of each of these prompts for each task, denoting these with a subscript S for sequence completion prompts and T for transducer prompts.
-Full listings of these prompts can be found in Appendix F.
-While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe our set of prompts captures common prompting strategies.
+Full listings of these prompts can be found in Appendix I.
+While no finite set of prompts will be fully sufficient to capture all possible model behavior, we believe maximizing over both Basic prompts allows both models that perform best at next-token-prediction and those that perform better in chain-of-thought reasoning to do their best.
+For each open weight model, we used a local VLLM (Kwon et al., 2023) server for evaluation and always evaluated on 1000 distinct DFAs.
+For GPT-4o and Claude, o3-mini, and gpt-5, we evaluated on 30 DFAs due to computation costs.
+(Due to greater interest in o3-mini’s performance on Red-GreenT, we used 100 to get a more precise estimate).
+For gpt-3.5 and gpt-4o-mini, we evaluated on 100 DFAs.
+All models were evaluated with temperature 0, except reasoning models o3-mini and gpt-5, which do not support a custom temperature.
 Model Size IT? Code? Sequence Completion SR Transducer TR
@@
 4-Gram 89.6 (88.7–90.4) 4 91.1 (90.6–91.6) 4
-3-Gram 87.0 (86.1–87.8) 5 87.0 (86.4–87.6) 16
-2-Gram 83.3 (82.2–84.2) 8 74.5 (73.6–75.3) 25
+3-Gram 87.0 (86.1–87.8) 5 87.0 (86.4–87.6) 19
+2-Gram 83.3 (82.2–84.2) 7 74.5 (73.6–75.3) 30
 Common-Suffix 84.7 (83.6–85.6) 6
-RandomS/NullT 53.3 (51.7–54.7) 26 68.9 (68.2–69.6) 26
-Open Source Completion
-llama3-8B 8.0B 73.8 (72.4–75.1) 18 87.5 (86.9–88.0) 14
-llama3-70B 70.6B 71.4 (70.0–72.7) 23 87.7 (87.2–88.3) 12
-llama3.1-8B-Instruct 8.0B 75.3 (74.0–76.6) 16 85.9 (85.3–86.5) 18
-mistral-nemo-minitron-8B 8.4B 78.7 (77.5–79.8) 12 88.6 (88.0–89.1) 5
-mistral-nemo-base-12B 12.2B 75.5 (74.3–76.6) 15 87.9 (87.4–88.4) 10
-mistral-nemo-instruct-12B 12.2B 72.2 (70.9–73.4) 22 88.0 (87.5–88.5) 8
-gemma-7b 8.5B 72.6 (71.3–73.7) 20 82.1 (81.4–82.7) 22
-falcon-7b 7.2B 69.0 (67.6–70.2) 24 84.9 (84.3–85.5) 20
-Open Source Code
-starcoder2-15b 16.0B 73.5 (72.0–74.7) 19 87.7 (85.8–89.5) 13
-codestral-22B 22.2B 78.0 (76.8–79.1) 13 86.6 (86.0–87.1) 17
-deepseek-coder-33b-instruct 33.3B 76.7 (75.3–77.8) 14 85.6 (85.0–86.2) 19
-qwen-2.5-coder-7B 7.6B 79.5 (78.4–80.5) 9 88.2 (87.6–88.7) 7
-qwen-2.5-coder-instruct-7B 7.6B 79.5 (78.3–80.5) 10 88.3 (87.8–88.8) 6
-qwen-2.5-coder-instruct-32B 32.8B 79.2 (78.0–80.3) 11 87.9 (87.4–88.4) 9
+RandomS/NullT 53.3 (51.7–54.7) 32 68.9 (68.2–69.6) 31
+Open Weight Completion
+llama3-8B 8.0B 73.8 (72.4–75.1) 22 87.5 (86.9–88.0) 18
+llama3-70B 70.6B 71.4 (70.0–72.7) 29 87.7 (87.2–88.3) 15
+llama3.1-8B-Instruct 8.0B 75.3 (74.0–76.6) 19 85.9 (85.3–86.5) 22
+llama3.1-8B 8.0B 75.2 (73.8–76.3) 20 88.0 (87.5–88.6) 10
+llama3.1-70B 70.0B 71.8 (70.4–73.1) 28 87.7 (87.2–88.2) 17
+qwen-2.5-7B 7.6B 73.5 (72.1–74.8) 24 88.7 (88.2–89.2) 5
+qwen-2.5-32B 32.5B 76.8 (75.5–78.0) 16 88.3 (87.8–88.8) 7
+mistral-nemo-minitron-8B 8.4B 78.7 (77.5–79.8) 13 88.6 (88.0–89.1) 6
+mistral-nemo-base-12B 12.2B 75.5 (74.3–76.6) 18 87.9 (87.4–88.4) 13
+mistral-nemo-instruct-12B 12.2B 72.2 (70.9–73.4) 27 88.0 (87.5–88.5) 11
+gemma-7b 8.5B 72.6 (71.3–73.7) 25 82.1 (81.4–82.7) 27
+falcon-7b 7.2B 69.0 (67.6–70.2) 30 84.9 (84.3–85.5) 24
+Open Weight Code
+starcoder2-15b 16.0B 73.5 (72.0–74.7) 23 87.7 (85.8–89.5) 16
+codestral-22B 22.2B 78.0 (76.8–79.1) 14 86.6 (86.0–87.1) 21
+deepseek-coder-33b-instruct 33.3B 76.7 (75.3–77.8) 17 85.6 (85.0–86.2) 23
+qwen-2.5-coder-7B 7.6B 79.5 (78.4–80.5) 10 88.2 (87.6–88.7) 9
+qwen-2.5-coder-instruct-7B 7.6B 79.5 (78.3–80.5) 11 88.3 (87.8–88.8) 8
+qwen-2.5-coder-instruct-32B 32.8B 79.2 (78.0–80.3) 12 87.9 (87.4–88.4) 12
 Proprietary
-gpt-3.5-instruct ? 67.3 (63.1–71.5) 25 87.8 (85.9–89.6) 11
-gpt-3.5-chat ? N/A 66.8 (63.4–69.8) 27
-gpt-4o-mini ? 72.4 (68.1–76.3) 21 79.8 (77.3–82.2) 23
-gpt-4o ? 74.4 (69.9–78.6) 17 83.7 (80.1–86.9) 21
-claude-3.5 ? 84.0 (79.3–88.4) 7 87.1 (83.9–90.2) 15
-o1-preview ? 76.5 (69.4–84.3) 24
-Table 1: Results for our experiments.
-We present model metadata alongside model results on both the Transducer and Sequence completion tasks.
-Each cell contains the mean performance across DFAs for the best-performing prompt (see Table 2 for details), with 95% confidence intervals of the mean in parentheses.
+gpt-3.5-instruct ? 67.3 (63.1–71.5) 31 87.8 (85.9–89.6) 14
+gpt-3.5-chat ? N/A 66.8 (63.4–69.8) 32
+gpt-4o-mini ? 72.4 (68.1–76.3) 26 79.8 (77.3–82.2) 28
+gpt-4o ? 74.8 (69.3–80.4) 21 83.7 (80.1–86.9) 25
+claude-3.5 ? 82.8 (77.5–87.5) 8 86.9 (83.3–90.0) 20
+o3-mini ? 81.1 (76.0–85.8) 9 74.7 (70.7–78.8) 29
+gpt-5 ? 77.9 (71.6–84.0) 15 83.6 (79.9–87.1) 26
+Table 1: Results for our experiments on both the Transducer and Sequence completion tasks.
+Each cell contains the mean performance across DFAs for the best-performing Basic prompt (see Table 2 for details), with 95% confidence intervals of the mean in parentheses.
 “N/A” is used whenever the model returned an invalid result at least 25% of the time.
-(IT = Instruction-Tuned, TR/SR = Transducer/Sequence Completion rank, the ordinal rank of the given model on the given task.)
+(IT = Instruction-Tuned, TR/SR = rank of the given model on Transducer/Sequence Completion.)
+Model Basic Basic-COT More-Expl DFA-COT Red-Green
+Sequence Completion
+gpt-4o-mini 72.4 (68.1–76.3) 60.0 (55.8–64.4) 70.5 (66.4–74.6) 58.0 (53.4–62.4) 59.1 (54.9–63.2)
+gpt-4o 72.1 (65.9–78.2) 74.8 (69.3–80.4) N/A 67.4 (60.8–73.8) 74.4 (69.9–78.6)
+claude-3.5 N/A 82.8 (77.5–87.5) N/A 84.0 (79.3–88.4) 80.0 (74.9–85.2)
+o3-mini N/A 81.1 (76.0–85.8) N/A 58.2 (49.6–66.8) 69.8 (64.4–75.0)
+gpt-5 77.9 (71.6–84.0) 75.1 (68.6–81.7) 68.9 (61.0–76.7) 66.0 (58.9–72.8) 87.5 (83.1–91.5)
+Transducer
+gpt-4o-mini 79.8 (77.3–82.2) 66.5 (64.3–68.7) 76.7 (74.2–79.3) 65.2 (63.1–67.4) 74.5 (72.0–77.0)
+gpt-4o 83.7 (80.1–86.9) 67.8 (62.4–73.2) 82.6 (79.1–85.9) 67.8 (63.1–72.3) 82.6 (78.8–86.3)
+claude-3.5 86.9 (83.3–90.0) 74.2 (70.0–78.3) 87.1 (83.9–90.2) 76.4 (72.9–79.9) 82.9 (78.9–86.9)
+o3-mini 72.8 (68.4–77.3) 74.7 (70.7–78.8) 74.7 (70.3–79.2) 86.1 (83.9–88.4) 92.4 (91.3–93.5)
+gpt-5 83.6 (79.1–87.7) 83.6 (79.9–87.1) 85.2 (81.0–88.8) 96.7 (95.3–98.0) 96.6 (95.4–97.8)
+Table 2: Results for models where we investigated multiple prompts (we only used Basic on other models).
+We bold the best prompt for each model.
+Non-COT prompts consistently work better for the Transducer task, with more mixed results on sequence completion.
 5 Results
 Main results for all tasks are presented in Table 1.
-For all LLMs, we ignore non-answers, i.e., if for a given DFA a model gets 25 correct answers, 1 incorrect answer, and responds with an unparseable result on 4, this counts as a 25/26, not a 25/29.
-We then report the mean across DFAs and 95% bootstrap confidence intervals.
+We ignore non-answers, i.e., if for a given DFA a model gets 25 correct answers, 1 incorrect answer, and responds with an unparseable result on 4, this counts as a 25/26, not a 25/29.
+We then aggregate across DFAs.
+All comparisons involving 4-Gram, 5-Gram, and 6-Gram to all other models are statistically significant (see Appendix H for details).
 5.1 Sequence Completion
-As seen in Table 1, this task is nearly always fully determined, that is, it can be solved with $\sim$100accuracy in theory, as demonstrated by BruteForceS results.
+As seen in Table 1, this task is nearly always fully determined, that is, it can be solved with $\sim$100% accuracy in theory, as demonstrated by BruteForceS results.
 Of course, BruteForceS is extremely computationally expensive, and, as such, we primarily focus on the $n$-GramS heuristics as our baselines.
@@
 As seen in Table 2, we find that giving the model the opportunity to logically reason about the prompt via chain-of-thought and present a conclusion has inconsistent results.
-Specifically, we find that BasicS is the best prompt for gpt-4o-mini, but not gpt-4o, where the best performing prompt is Red-GreenS.
-We find that claude-3.5 is entirely unable to follow the sequence completion prompts BasicS and More-ExplS, and performs best at the COTS prompt.
-Additionally, we find that in this task, code-specific open-source models tend to perform better than sequence completion models, suggesting some generalized ability to produce strings from novel languages demonstrated by example.
-Overall, the relative performances of LLMs and prompts comport somewhat well to heuristics on which models and prompting strategies should work best.
+Specifically, we find that for gpt-4o-mini, immediately predicting a next token seems to be better, while for gpt-4o and gpt-5 there is no large effect.
+claude-3.5 and o3-mini are unable to answer the BasicS prompt at all, but outperform other proprietary models when using Basic-COTS.
+In this task, revealing the problem structure appears to not have a massive effect on performance, with gpt-5 being the only model to incorporate this information into a statistically significantly improved performance, and even then only in the Red-GreenS word problem prompt (still underperforming 4-GramS).
+Additionally, we find that in this task, code-specific open-weight models tend to perform better than sequence completion models, suggesting some generalized ability to produce strings from novel languages demonstrated by example.
+Overall, the relative performances of LLMs and prompts generally comport to heuristics on which models and prompting strategies should work best (with the notable exception of gpt-5).
 Nonetheless, LLMs underperform simple $n$-Gram heuristics.
 One potential problem with using this task for cross-model comparisons is the relevance of tokenization.
-Unfortunately, we found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix E for details.
-Model Basic More-Expl COT Red-Green
-Sequence Completion
-gpt-4o-mini 72.4 (68.1–76.3) 70.5 (66.4–74.6) 58.0 (53.4–62.4) 59.1 (54.9–63.2)
-gpt-4o 72.1 (65.9–78.2) N/A 67.4 (60.8–73.8) 74.4 (69.9–78.6)
-claude-3.5 N/A N/A 84.0 (79.3–88.4) 80.0 (74.9–85.2)
-Transducer
-gpt-4o-mini 79.8 (77.3–82.2) 76.7 (74.2–79.3) 65.2 (63.1–67.4) 74.5 (72.0–77.0)
-gpt-4o 83.7 (80.1–86.9) 82.6 (79.1–85.9) 67.8 (63.1–72.3) 82.6 (78.8–86.3)
-claude-3.5 86.9 (83.3–90.0) 87.1 (83.9–90.2) 76.4 (72.9–79.9) 82.9 (78.9–86.9)
-Table 2: Results for models where we investigated multiple prompts (we only used Basic on other models).
-We bold the best prompt for each model.
-Non-COT prompts consistently work better for the Transducer task, with more mixed results on sequence completion.
+We found that forcing uniform tokenization by using commas in the prompt uniformly reduced accuracy, see Appendix F.1 for details; we confirmed that the LLMs we investigate could reason about strings without commas, see Appendix F.2 for details.
 5.2 Transducer
-Unlike sequence completion, this task is generally not fully determined, with the $\textsc{BruteForce}_{T}$ baseline only achieving 96.4% accuracy.
-Comparisons are still be valid as all models see the same fraction of unsolvable instances.
+Unlike sequence completion, this task is not fully determined, with $\textsc{BruteForce}_{T}$ getting 96.4% accuracy.
+Comparisons are still valid as all models see the same fraction of unsolvable instances.
 We find that in general all LLMs underperform a 4-GramT model, demonstrating that they are unable to adequately solve this task.
-The relative performance of the models also does not correspond to their overall scale, with open source LLama-3 and Mistral Nemo 8B parameter models outperforming Claude and GPT-4o.
-Even within a model class we find no clear pattern: GPT-4o and o1-preview [^4] The particularly poor performance of o1-preview may be due to the model not supporting temperature 0.
-In other experiments, we found that GPT models with temperature 1 tended to perform poorly.
-See Appendix B for more details.
-are outperformed by GPT 3.5, Llama 3-70B has similar performance to Llama 3-8B, and the Mistral Nemo 12B models perform similarly to Nemo Minitron 8B.
+The relative performance of the models also does not correspond to their overall scale, with open weight LLama-3 and Mistral Nemo 8B parameter models outperforming much larger proprietary models.
+Even within a model class we find no clear pattern: all other GPT models are outperformed by GPT 3.5, Llama 3-70B has similar performance to Llama 3-8B, and the Mistral Nemo 12B models perform similarly to Nemo Minitron 8B.
 Coding models also demonstrate no advantage on this task.
 The generally lower performance of chat-oriented models suggests this task is better suited to non-chat models.
-To investigate that this is not specific to the Basic prompt, we investigate other prompts for chat models.
-As seen in Table 2, our chain-of-though and word problem prompts, which attempt to leverage the full reasoning capabilities of chat models, also fare poorly, performing similarly or worse to the Basic prompt on the Transducer task in all cases.
-We conclude that LLMs are unable to perform well on the DFA transducer inference task.
+More specifically, as seen in Table 2, our Basic-COTT prompt results in underperformance by all non-reasoning models, suggesting that models are generally most able to solve this task when it is a simple next-token-prediction task.
+Providing the problem structure also does not help non-reasoning models improve substantially, but does allow o3-mini to perform well, and gpt-5 to completely solve the task (achieving parity with BruteForce), demonstrating that reasoning models’ underperformance at the original ICL language reasoning task is not due to the underlying difficulty of the task itself.
+We conclude that LLM ICL is unable to perform well at language inference.
 This failure cannot be attributed to a lack of world modeling, as $n$-GramT models do not construct world models.
-Instead, it seems the LLMs are unable to detect patterns when those patterns are drawn from an unfamiliar source, even a relatively simple one.
+Instead, it seems the LLMs are unable to detect patterns when those patterns are drawn from an unfamiliar and unknown source, even a relatively simple one.
 5.3 Comparison of Benchmarks
 Figure 3: Transducer and sequence completion results plotted against each other.
-Points represent the mean over several DFAs and intervals represent 95% confidence intervals.
-Points are colored by model type, with the best and worst model by each metric in each category labeled, as well as all baseline and proprietary models.
+Points are the mean over several DFAs, with 95% confidence intervals.
+Points are colored by model type, with the best and worst model by each metric in each category labeled, as well as all baseline & proprietary models.
 Figure 3 displays the relationship between model performance on the Sequence Completion and Transducer benchmarks.
 While at a high level, there is a positive correlation between the two, there are a few notable differences.
-For one, the Code models perform notably better than other open source models on Sequence Completion, but not on Transducer.
-Additionally, on Transducer, a ceiling on performance is observed, where LLMs cluster together between 3-GramT and 4-GramT performance; this clustering does not appear on the Sequence Completion benchmark.
-5.4 Case Study: Sum Modulo 3 DFA
-Correct Incorrect
-Total 100% 100%
-a is no-op 70% 73%
-1b and 1c lead to 0 47% 57%
-2-periodic 30% 47%
-3-periodic 13% 13%
-2 red rooms 7% 10%
-Figure 4: Results on Sum Modulo 3 DFA.
-(a) MB=mistral-nemo-minitron-8B/BasicT, CR=claude-3.5/Red-GreenT.
-Venn diagram of errors (out of 1000).
-Labeled percentages are accuracies.
-(b) Results of qualitative analysis, out of 30 in both cases.
-We investigate the transducer task on the DFA depicted in Figure 1.
-This DFA can be interpreted as an arithmetic check, where a represents 0, b represents 1, and c represents 2, and the DFA accepts strings whose sum is equal to 0 modulo 3.
-For this case study, we focus on the model/prompt combinations MB and CR, defined as
-  • MB: mistral-nemo-minitron-8B/BasicT. Selected as it is the best performing combination overall.
-  • CR: claude-3.5/Red-GreenT. Selected as it is the best performing combination that provides an explanation (needed later for our qualitative analysis)
-Figure 4a depicts the number of errors each model receives on 1000 instances of the transducer task for this DFA.
-Nearly all errors made by the 6-GramT model were also made by at least one LLM, while the two LLMs often made unique errors.
-While this task is better-known than most DFAs, all 3 models perform worse on this DFA than average.
-We also performed a qualitative analysis, investigating CR’s outputs on the Red-GreenT prompt to see what kind of reasoning it is using; specifically we sampled 30 examples where it had the correct answer, and 30 examples where it had the incorrect answer but the 6-GramT model had the correct answer.
-Results of this analysis can be found in Figure 4b.
-We find that, in general, CR is following a 3-Gram approach, learning rules relating to the conditions under which the previous output and symbol can be used to predict the next output.
-Specifically, it is able to learn that a does not change the output, and that b and c will lead a 1 state to a 0 state.
-These results comport with the overall finding of Table 1, where we found that 3-GramT was the largest $n$-GramT that any LLM outperformed, as well as our finding that LLM performance decreases for tasks that are not solvable by $n$-Grams; see Appendix C for details.
-[…]
-Despite transformers’ high computational capacity, without the ability to pattern match to existing problems, Claude uses an unsophisticated and ineffectual approach.
+For one, the Code models perform notably better than other open weight models on Sequence Completion, but not on Transducer.
+Additionally, on Transducer, a ceiling on performance is observed, where non-reasoning LLMs cluster together between 3-GramT and 4-GramT performance; this clustering does not appear on the Sequence Completion benchmark.
 6 Conclusion
-Our findings highlight significant weaknesses in large language models’ ability to generalize to entirely novel language reasoning problems, even simple ones solely involving next-token prediction on basic languages recognized by 3-state DFAs.
+Our findings highlight significant weaknesses in large language models’ ability to in-context-learn entirely novel language reasoning problems, even simple ones solely involving next-token prediction on basic languages recognized by 3-state DFAs.
 These results, combined with that of previous work demonstrating that large language models can quite accurately perform a variety of language tasks, suggests that LLMs solve language problems via a mechanism distinct from general language reasoning ability.
@@
 We believe our results suggest that LLMs have learned individual models of particular languages, but not a general theory of language.
-Interestingly, in our transducer experiments, LLMs consistently perform better by directly predicting the next token than by explicitly reasoning through the problem.
+Interestingly, in our transducer experiments, directly predicting the next token rather than explicitly reasoning through the problem works better, except for reasoning models where they perform similarly.
 While our conclusions are limited by the finite nature of our prompt set, this suggests that they do, in fact, possess some latent understanding of language, but this understanding is inferior to basic n-gram models for $n>3$.
-Many potential foundation model applications involve tasks that are not expressed in familiar human languages or pre-existing programming languages.
-More specifically, in tasks where there is a need to produce an output in a precise, atypical, format, we should be skeptical of the ability of LLMs to in-context-learn this format.
-For these tasks, it may be prudent to seek a new approach.
+One potential goal for foundation models is to replace all machine learning with ICL. Our results suggest that current models are not progressing towards this goal.
 Impact Statement
@@
 If every transducer trace ends with a 0 or every trace ends with a 1, we reject the DFA and resample.
-Appendix B Results of o1-preview
-We evaluated o1-preview on 10 DFAs, using 30 problem instances per DFA of the Transducer task, as in other Transducer experiments, and the $\textsc{Basic}_{T}$ prompt, as this is the most neutral prompt.
-Table 3 displays results on each DFA.
-Overall, while these results are not on a particularly large sample, they fairly definitively demonstrate that o1-preview does not achieve strikingly good performance on this task.
-DFA o1-preview gpt-4o 6-Gram
-1 25/30 27/30 26/30
-2 23/29 24/30 25/30
-3 19/30 23/30 28/30
-4 22/30 23/30 28/30
-5 29/29 30/30 30/30
-6 19/30 24/30 30/30
-7 17/29 23/30 25/30
-8 23/30 25/30 26/30
-9 21/30 28/30 30/30
-10 29/30 29/30 30/30
-Table 3: Results on each DFA.
-We find that in 7 cases, o1-preview underperforms gpt-4o, in 2 cases it gets the same number of instances wrong but provides a non-answer on an additional instance, and in 1 case it ties gpt-4o.
-In no cases does it outperform.
-Appendix C Transducer results by difficulty class
+Appendix B Transducer results by difficulty class
 Figure 4: Transducer results by difficulty class.
@@
 Additionally, we find that the best models continue to perform similarly to 4-Gram for tasks that 4-Gram does not perfectly solve.
-Appendix D More details on Sum Modulo 3 DFA case study
-Figure 6 depicts the results of the Sum Modulo 3 experiment, but filtered for two conditions.
-In the (a) condition, the trace ends in such a way that a 3-Gram model would be able to determine the output, and the (b) condition is the complement.
+Appendix C Case Study: Sum Modulo 3 DFA
+Correct Incorrect
+Total 100% 100%
+a is no-op 70% 73%
+1b and 1c lead to 0 47% 57%
+2-periodic 30% 47%
+3-periodic 13% 13%
+2 red rooms 7% 10%
+Figure 5: Results on Sum Modulo 3 DFA.
+(a) MB=mistral-nemo-minitron-8B/BasicT, CR=claude-3.5/Red-GreenT.
+Venn diagram of errors (out of 1000).
+Labeled percentages are accuracies.
+(b) Results of qualitative analysis, out of 30 in both cases.
+We investigate the transducer task on the DFA depicted in Figure 1.
+This DFA can be interpreted as an arithmetic check, where a represents 0, b represents 1, and c represents 2, and the DFA accepts strings whose sum is equal to 0 modulo 3.
+For this case study, we focus on the model/prompt combinations MB (mistral-nemo-minitron-8B/BasicT: the best performing non-reasoning model combination overall) and CR (claude-3.5/Red-GreenT: the best performing non-reasoning model combination that provides an explanation, needed later for our qualitative analysis).
+Figure 5a depicts the number of errors each model receives on 1000 instances of the transducer task for this DFA.
+Nearly all errors made by the 6-GramT model were also made by at least one LLM, while the two LLMs often made unique errors.
+While this task is better-known than most DFAs, all 3 models perform worse on this DFA than average.
+We also performed a qualitative analysis, investigating CR’s outputs on the Red-GreenT prompt to see what kind of reasoning it is using; specifically we sampled 30 examples where it had the correct answer, and 30 examples where it had the incorrect answer but the 6-GramT model had the correct answer.
+Results of this analysis can be found in Figure 5b.
+We find that, in general, CR is following a 3-Gram approach, learning rules relating to the conditions under which the previous output and symbol can be used to predict the next output.
+Specifically, it is able to learn that a does not change the output, and that b and c will lead a 1 state to a 0 state.
+These results comport with the overall finding of Table 1, where we found that 3-GramT was the largest $n$-GramT that any non-reasoning LLM outperformed, as well as our finding that LLM performance decreases for tasks that are not solvable by $n$-Grams; see Appendix B for details.
+The model also attempts to identify periodic patterns, but identifies period-2 patterns more than period-3 patterns, despite knowing that there are three “rooms” (states).
+At no point in any of the 60 reasoning traces analyzed does it realize that this is a version of the Sum Modulo 3 DFA [^5] In fact, in none of the 1000 traces do the substrings “sum” or “mod” appear, except as a part of “assuming”, but it does show some glimmers of world modeling: in a few cases it correctly determines that there are two red rooms; however, this does not lead to further discoveries.
+It is not superior reasoning that leads to correct solutions, rather the correct examples are more likely to be ones that a 3-Gram model would infer correctly, i.e., those traces ending in a, 1b, or 1c, which occur cumulatively in $\frac{5}{9}$ of cases [^6] On the $\sim$$\frac{5}{9}$ of examples following this pattern, CR achieves 93.5%, to the 6-GramT’s 97.3%, and on the remaining $\sim$$\frac{4}{9}$, it achieves 43.8%, to the 6-GramT’s 60.7%.
+Detailed Venn diagrams on these conditions can be found in Figure 6..
+Despite transformers’ high computational capacity, without the ability to pattern match to existing problems, Claude uses an unsophisticated and ineffectual approach.
 Figure 6: Results on Sum Modulo 3 DFA under trivial / nontrivial conditions.
@@
 Here, errors are more symmetric, with more 6-GramT errors that are not accounted for by either or both model, indicating that a larger fraction of both successes and failures in this condition are down to random chance.
+Appendix D Increased number of examples
+Figure 7: Accuracy by number of examples, as we vary the number from 30 to 600 (note the log scale).
+$\infty$-Gram is a model that finds the longest sequence that matches the ending of the sequence and copies the following token.
+We varied the number of examples parameter from 30 to 600, investigating specifically mistral-nemo-minitron-8B as it is a better performing model on the transducer task.
+We find that for the Sequence Completion task, the $n$-Gram models do not generally improve as the number of examples increases, while the LLM does; however, the LLM remains below 4-Gram performance at all points.
+In the Transducer experiment, both the n-Gram models and the LLM improve, with the LLM crossing 4-Gram and 5-Gram performance and eventually roughly matching 6-Gram performance.
+However, there is now an increased gap between 6-Gram and 7-Gram that did not exist at 30 (we did not include 7-Gram or above in the main table for this reason).
+Overall, it is possible that the larger number of instances of the “correct n-Gram” appearing (i.e., the suffix of the prompt followed by the correct answer) causes the model to be better at producing predictions.
+As the LLM at no point outperforms the best n-Gram, we do not believe that this alters our overall conclusions.
+Appendix E Nonzero temperature
+Model Prompt Zero Temp Nonzero Temp Difference
+Sequence Completion
+mistral-nemo-minitron-8B Basic 78.70% (77.49% – 79.79%) 77.67% (76.51% – 78.76%) -1.04% (-1.49% – -0.63%)
+claude-3.5 COT 84.00% (79.33% – 88.44%) 84.22% (79.56% – 89.00%) 0.22% (-2.11% – 2.33%)
+claude-3.5 Red-Green 80.00% (74.89% – 85.22%) 80.78% (75.00% – 86.11%) 0.78% (-2.11% – 3.44%)
+Transducer
+mistral-nemo-minitron-8B Basic 88.56% (88.05% – 89.08%) 88.17% (87.64% – 88.68%) -0.39% (-0.58% – -0.22%)
+claude-3.5 Basic 86.89% (83.33% – 90.00%) 87.00% (83.33% – 90.11%) 0.11% (-0.67% – 0.89%)
+claude-3.5 More-Expl 87.11% (83.88% – 90.22%) 86.89% (83.22% – 90.11%) -0.22% (-1.11% – 0.67%)
+claude-3.5 COT 76.44% (72.89% – 79.89%) 78.11% (74.66% – 81.33%) 1.67% (-0.56% – 3.78%)
+claude-3.5 Red-Green 82.89% (78.89% – 86.89%) 82.78% (78.89% – 86.33%) -0.11% (-2.00% – 1.78%)
+Table 3: Results varying temperature.
+Second column is a temperature of 0.1, third column is differences.
+In all columns, we annotate a 95% confidence interval, using paired differences for the third column.
+In order to determine whether a small nonzero temperature might lead to better results, we investigated using a temperature of 0.1 for mistral-nemo-minitron-8B and claude-3.5 (these two chosen for the reasons described in Section C, notably o3-mini/gpt-5 already are using nonzero temperatures).
+We find that a temperature of 0.1 does not significantly change performance, resulting in no significant change for any model/task/prompt combination.
+The largest improvement we observe is claude-3.5 on COT on the Transducer task, which gains 1.7%, from 76.4% to 78.1% (still not enough to make it the best prompt for claude-3.5).
+Appendix F Tokenization
 Model BasicS Basic-CommasS
@@
 codestral-22B 78.0 (76.8–79.1) 59.0 (57.5–60.3)
+qwen-2.5-32B 76.8 (75.5–78.0) 60.3 (58.9–61.6)
 deepseek-coder-33b-instruct 76.7 (75.3–77.8) 54.9 (53.0–56.8)
@@
 llama3.1-8B-Instruct 75.3 (74.0–76.6) 56.3 (54.4–58.1)
+llama3.1-8B 75.2 (73.8–76.3) 61.1 (59.8–62.5)
 llama3-8B 73.8 (72.4–75.1) 61.5 (60.2–62.9)
 starcoder2-15b 73.5 (72.0–74.7) 58.2 (56.7–59.8)
+qwen-2.5-7B 73.5 (72.1–74.8) 57.0 (55.5–58.5)
 gemma-7b 72.6 (71.3–73.7) 54.0 (51.9–56.0)
@@
 gpt-4o 72.1 (65.9–78.2) 66.8 (58.5–74.8)
+llama3.1-70B 71.8 (70.4–73.1) 57.7 (56.1–59.2)
 llama3-70B 71.4 (70.0–72.7) 56.4 (54.7–58.0)
@@
 gpt-3.5-instruct 67.3 (63.1–71.5) 52.3 (46.5–57.9)
+o3-mini N/A N/A
 claude-3.5 N/A N/A
@@
 Table 4: Results on Sequence Completion Task. We compare BasicS to the comma-variant Basic-CommasS.
-Appendix E Sequence Completion task prompt with Commas
+F.1 Sequence Completion task prompt with Commas
 To avoid tokenization differences with models, we also investigate a version of our Sequence Completion prompt that uses spaces and commas between the elements of the sequence.
@@
 All are worse with commas than without.
-Appendix F Prompt Listings
+F.2 Directly confirming models can read sequences of letters
+To fully exclude the possibility that models are unable to read sequences of letters, we perform the following experiment: we take the regular expression ab(abc)+ and directly provide it to the model, then ask the model to test it on a string (one string provided per query), using the prompt I will give you a string.
+Tell me whether it matches the following regular expression: ’^ab(abc)+\$’ (without quotes).
+Just answer YES or NO.
+on one line, followed by a string on the next.
+We sample 100 random strings of length 2-17, via the following procedure (1) sample a random valid string uniformly (2) with 50% probability, randomly mutate one of the elements of the string to a different character.
+All strings are thus either correct or near-correct.
+We also add Answer (YES or NO): on a third line for non-chat models to encourage them to provide a response rather than another string.
+Model Accuracy Non-response
+llama3-8B 66%
+llama3-70B 80% 26%
+llama3.1-8B-Instruct 83%
+llama3.1-8B 89%
+llama3.1-70B 86% 36%
+qwen-2.5-7B 62%
+qwen-2.5-32B 100%
+mistral-nemo-instruct-12B 99%
+gemma-7b 67% 15%
+starcoder2-15b 62%
+codestral-22B 95%
+qwen-2.5-coder-7B 100%
+qwen-2.5-coder-instruct-7B 56%
+qwen-2.5-coder-instruct-32B 96%
+gpt-3.5-instruct 67%
+gpt-3.5-chat 82%
+gpt-4o-mini 100%
+gpt-4o 100%
+claude-3.5 100%
+o3-mini 100%
+gpt-5 100%
+Table 5: Results on Regex task.
+As in the rest of this paper, accuracies are computed ignoring non-response.
+Models mistral-nemo-minitron-8B, mistral-nemo-base-12B, falcon-7b, deepseek-coder-33b-instruct have non-response rates over 98%.
+Results for this are presented in Table 5.
+While not all models perform well at this task, performances follow roughly what one might expect from standard benchmarks (rather than the Sequence Completion results in the paper).
+Frontier proprietary models perform extremely well, larger open weight models tend to perform well, and smaller open weight models are more hit-and-miss (some performing well, some poorly).
+The proprietary models newer than the GPT-3 series get 100% accuracy on this task, while performance is lower for the open weight models.
+There is no obvious relationship between models that perform well at this particular task and models that perform well at Sequence Completion, indicating that to whatever degree models are performing poorly at this test task, it is not because they are unable to process the string.
+Notably, o3-mini, gpt-4o, and gpt-4o-mini all perform fairly poorly at the Sequence Completion task, placing below median, but all achieve 100% on this task.
+Open weight models are more mixed, with larger ones tending to do nearly as well as newer proprietary models (notable exception being llama3-70B), but smaller ones occasionally performing well and occasionally performing poorly.
+Appendix G Model non-answers
+Model Basic Basic-COT More-Expl DFA-COT Red-Green
+Sequence Completion
+llama3-8B 0.0 (0.0–0.0)
+llama3-70B 0.0 (0.0–0.0)
+llama3.1-8B-Instruct 0.0 (0.0–0.0)
+llama3.1-8B 0.0 (0.0–0.0)
+llama3.1-70B 0.0 (0.0–0.0)
+mistral-nemo-minitron-8B 0.0 (0.0–0.0)
+mistral-nemo-base-12B 0.0 (0.0–0.0)
+mistral-nemo-instruct-12B 0.0 (0.0–0.0)
+gemma-7b 0.0 (0.0–0.0)
+falcon-7b 0.0 (0.0–0.0)
+starcoder2-15b 0.0 (0.0–0.0)
+codestral-22B 0.0 (0.0–0.0)
+deepseek-coder-33b-instruct 0.0 (0.0–0.0)
+qwen-2.5-coder-instruct-7B 0.0 (0.0–0.0)
+qwen-2.5-7B 0.0 (0.0–0.0)
+qwen-2.5-32B 0.0 (0.0–0.0)
+qwen-2.5-coder-instruct-32B 0.0 (0.0–0.0)
+qwen-2.5-coder-7B 0.0 (0.0–0.0)
+gpt-3.5-instruct 2.5 (1.9–3.0)
+gpt-3.5-chat 99.9 (99.7–100.0)
+gpt-4o-mini 0.0 (0.0–0.0) 0.3 (0.2–0.5) 0.0 (0.0–0.0) 1.0 (0.6–1.4) 0.2 (0.1–0.4)
+gpt-4o 4.4 (3.3–5.7) 1.9 (1.1–2.8) 100.0 (100.0–100.0) 5.0 (3.8–6.2) 8.4 (6.6–10.2)
+claude-3.5 99.7 (99.2–100.0) 0.1 (0.0–0.3) 97.8 (96.9–98.6) 0.0 (0.0–0.0) 0.0 (0.0–0.0)
+o3-mini 80.2 (78.0–82.3) 3.8 (2.6–5.2) 91.6 (89.9–93.2) 5.7 (4.1–7.2) 0.4 (0.0–1.0)
+gpt-5 0.0 (0.0–0.0) 0.0 (0.0–0.0) 8.2 (4.7–12.1) 5.2 (3.1–7.4) 0.4 (0.1–0.9)
+Transducer
+llama3-8B 0.0 (0.0–0.0)
+llama3-70B 0.0 (0.0–0.0)
+llama3.1-8B-Instruct 0.0 (0.0–0.0)
+llama3.1-70B 0.0 (0.0–0.0)
+llama3.1-8B 0.0 (0.0–0.0)
+starcoder2-15b 0.0 (0.0–0.0)
+codestral-22B 0.0 (0.0–0.0)
+deepseek-coder-33b-instruct 0.0 (0.0–0.0)
+qwen-2.5-coder-7B 0.0 (0.0–0.0)
+qwen-2.5-coder-instruct-7B 0.0 (0.0–0.0)
+qwen-2.5-7B 0.0 (0.0–0.0)
+qwen-2.5-32B 0.0 (0.0–0.0)
+qwen-2.5-coder-instruct-32B 0.0 (0.0–0.0)
+mistral-nemo-minitron-8B 0.0 (0.0–0.0)
+mistral-nemo-base-12B 0.0 (0.0–0.0)
+mistral-nemo-instruct-12B 0.0 (0.0–0.0)
+gemma-7b 0.0 (0.0–0.0)
+falcon-7b 0.0 (0.0–0.0)
+gpt-3.5-instruct 0.0 (0.0–0.1)
+gpt-3.5-chat 0.1 (0.0–0.3)
+gpt-4o-mini 1.8 (1.3–2.3) 0.0 (0.0–0.0) 5.8 (4.8–6.9) 0.0 (0.0–0.0) 0.7 (0.4–1.0)
+gpt-4o 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0)
+claude-3.5 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0)
+o3-mini 0.1 (0.0–0.3) 0.0 (0.0–0.0) 0.4 (0.1–0.9) 0.0 (0.0–0.0) 0.0 (0.0–0.1)
+gpt-5 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.0 (0.0–0.0) 0.7 (0.1–1.4)
+Table 6: Model non-answers, as a percentage of all prompt responses.
+A non-response is not included in accuracy computations for Table 1 or Table 2, but whenever it rises above 25%, N/A is placed in those tables.
+Table 6 depicts the percentage of model non-answers by model and prompt. In general, this distribution is highly bimodal, with values always being either below 9% or above 97%.
+The only prompt-vs-prompt orderings that are changed by scoring non-answers as 0 are that, on Sequence Completion, BasicS rises above Red-GreenS for gpt-4o, making it the best prompt; and that on Transducer, Red-GreenT for gpt-4o-mini rises above More-ExplT (though still behind BasicT.
+The qualitative conclusions about next token prediction vs chain of thought results remain the same.
+The only change to relative model ordering is that on Sequence Completion, gpt-4o drops 8 ranks, from 17th place to 25nd place, being passed by several open weight models, gpt-4o-mini, and o3-mini.
+No change occurs on the transducer results.
+Qualiative conclusions about model ordering remain the same.
+Appendix H Significance
+Figure 8: Significance of comparisons between rows of Table 1. Results in blue and orange are significant, results in pink are not.
+Figure 8 shows which comparisons between rows of Table 1 are significant. Significance computations are performed by running a 2-tailed bootstrap significance test on paired (by DFAs) differences.
+Appendix I Prompt Listings
 Prompt $T$ $S$
@@
 Output only the necessary suffix to complete the final string, and nothing else.
+<EXAMPLES> <PREFIX>
+Basic-COT You are a sequence completion model.
+Reason step by step, and then output the next output integer using <answer> tags, like <answer>0</answer>.
+Input sequence: <TRANSDUCER PREFIX> Output sequence: The following strings come from an alien language that follows a simple grammar.
+Infer the alien grammar using the example strings.
+Then, add a suffix to the final string using between 1 and 5 characters such that the full string follows the grammar.
+Reason step by step, and then output the next necessary suffix for this final string, <answer> tags, like <answer>ab</answer>.
 <EXAMPLES> <PREFIX>
@@
 <EXAMPLES> <PREFIX>
-COT A DFA is a finite-state machine that accepts or rejects a given string of symbols, by running through a n-state sequence uniquely determined by the string.
+DFA-COT A DFA is a finite-state machine that accepts or rejects a given string of symbols, by running through a n-state sequence uniquely determined by the string.
 I have a 3-state DFA model that outputs either 0 or 1 after each element I input.
@@
 Table 7: Shortened summary of each prompt
-F.1 Summaries
+I.1 Summaries
 Table 7 contains summaries of each prompt.
-F.2 Full example listings
-F.2.1 BasicT
+I.2 Full example listings
+I.2.1 BasicT
 You are a sequence completion model. Output the next element of the sequence, and nothing else.
 a, 1, b, 1, a, 1, b, 1, b, 1, c, 0, a, 1, c, 1, a, 1, a, 1, a, 1, c, 1, b, 1, c, 0, c, 1, a, 1, b, 1, b, 1, b, 1, b, 1, a, 1, b, 1, a, 1, a, 1, b, 1, c, 0, a, 1, c, 1, a, 1, b,
-F.2.2 More-ExplT
+I.2.2 Basic-COTT
+You are a sequence completion model. Reason step by step, and then output the next output integer using <answer> tags, like <answer>0</answer>.
+Input sequence: a, b, a, b, b, c, a, c, a, a, a, c, b, c, c, a, b, b, b, b, a, b, a, a, b, c, a, c, a, b
+Output sequence: 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1,
+I.2.3 More-ExplT
 You are a sequence completion model.
@@
 a, 1, b, 1, a, 1, b, 1, b, 1, c, 0, a, 1, c, 1, a, 1, a, 1, a, 1, c, 1, b, 1, c, 0, c, 1, a, 1, b, 1, b, 1, b, 1, b, 1, a, 1, b, 1, a, 1, a, 1, b, 1, c, 0, a, 1, c, 1, a, 1, b,
-F.2.3 COTT
+I.2.4 DFA-COTT
 A DFA is a finite-state machine that accepts or rejects a given string of symbols, by running through a n-state sequence uniquely determined by the string.
@@
 Output sequence: 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 0, 1, 1, 1,
-F.2.4 Red-GreenT
+I.2.5 Red-GreenT
 You are in a house of rooms and portals.
@@
 Then, you walk through a portal labeled "B" and end up in a ...
-F.2.5 BasicS
+I.2.6 BasicS
 The following strings come from an alien language that follows a simple grammar.
@@
 caacb
-F.2.6 Basic-CommasS
+I.2.7 Basic-COTS
+The following strings come from an alien language that follows a simple grammar.
+Infer the alien grammar using the example strings.
+Then, add a suffix to the final string using between 1 and 5 characters such that the full string follows the grammar.
+Reason step by step, and then output the next necessary suffix for this final string, <answer> tags, like <answer>ab</answer>.
+cbcbabbcca
+abcaaacbaa
+aabccbabbb
+bbbccbbbca
+aababaccba
+aaaacbacac
+baacbccbaa
+cbbaacabcc
+baabaacaab
+bbbbbcacab
+acaabcbbba
+acaacbccac
+cacbabcbba
+abcbcbcbcc
+ccaccccaba
+bcbcabbcca
+baabacabca
+caababacac
+bacacaccaa
+bcacbbbbca
+bcbbbcaccc
+ccabbcccbb
+bccbcabbca
+baacbabcbc
+ccacabccab
+caacbcaaab
+cacbaaccac
+aaccbcaabb
+abacabcaab
+bacbcbcaca
+caacb
+I.2.8 Basic-CommasS
 The following strings come from an alien language that follows a simple grammar.
@@
 c, a, a, c, b,
-F.2.7 More-ExplS
+I.2.9 More-ExplS
 I have a 3-state DFA model that outputs either 0 or 1 after each element I input.
@@
 caacb
-F.2.8 COTS
+I.2.10 DFA-COTS
 I have a 3-state DFA model that outputs either 0 or 1 after each element I input.
@@
 caacb
-F.2.9 Red-GreenS
+I.2.11 Red-GreenS
 You are outside a house of rooms and portals.
@@
 caacb
+Appendix J Compute Usage
+The experiments in this paper on proprietary models had the following (approximate) costs.
+  • gpt-5: \$920
+  • o3-mini: \$430
+  • 4o: somewhere between \$100 and \$200
+  • 4o-mini: somewhere between \$50 and \$150
+  • claude-3.5: \$80
+The open weight experiments took a cumulative 10-50 GPU-hours on NVIDIA RTX 6000 Ada Generation GPUs, some models requireed the use of 4 in parallel.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: We claim that the paper contributes a domain that demonstrates that novelty alone can cause LLMs to perform poorly.
+and in the paper we provide such a domain and comprehensive set of experiments to this effect.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We maintain a fairly narrow scope for the paper in terms of what we are trying to prove. We discuss limitations of our prompt set, which are the main direct limitations.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: we advance no novel formal theoretical claims in this work.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: we believe the results of this paper are as reproducible as is possible.
+A code repository contains the exact code used to run these experiments, which rely on data generated in that repository.
+The only significant concern for reproducibility is the availability of certain proprietary models.
+This issue is unfortunately unavoidable if we are to engage and evaluate proprietary models, so we do not view it as a significant limitation.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: A zip file has been attached containing the code.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: All information necessary to run the experiments is present in the paper. Many of the details are in Appendix A.1.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: 95% CIs are provided everywhere, and our main results (4-Gram through 6-Gram vs everything else) are all significant, as declared in the results section.
+For more details, see Appedix H.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: See Section J. o1-preview did not go into the paper as it was superceded by o3-mini, which performed better, cost an additional \$180.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: This experiment does not intersect any elements of the code of codnudct, except as discussed in the impact statement.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Yes, see post-conclusion impact statement.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: No models are being released as part of this paper. The dataset that is being released has no potential for problematic use, and would be fairly easy for someone to create anyway.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We cite every model we use, as well as VLLM. No other artifacts are used.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: no new assets are released.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: No crowdsourcing or human subjects were included in this paper.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: No crowdsourcing or human subjects were included in this paper.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79. Justification: LLM evaluation is the purpose of this paper, and as such LLMs were used throughout the experiments. Their use is precisely described.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.